### v2

# Motivic interpretations for iterated integrals on some specific algebraic curves

Motivic interpretations for iterated integrals on some specific algebraic curves
Eisuke Otsuka Eisuke Otsuka
Mathematical Institute, Tohoku University.
6-3, Aoba, Aramaki, Aoba-ku, Sendai, 980-8578, JAPAN
eisuke.otsuka.p3@dc.tohoku.ac.jp
Аннотация.
Multiple zeta values (MZVs for short) can be represented as iterated integrals of $\mathbb{Q}$-rational algebraic differential forms on $\mathbb{P}^{1}(\mathbb{C})\setminus\{0,1,\infty\}$. This interpretation allows us to consider MZVs geometrically, and this is one of the motivations for Deligne–Goncharov, Terasoma et al. to give motivic interpretations of MZVs by using the theory of mixed Tate motives and the motivic fundamental groups.
In this paper, we consider the iterated integrals on some rational curves over $\mathbb{Q}$ and study their arithmetic properties. They are an extension of MZVs and also include some other known special values such as multiple $\widetilde{T}$-values. Furthermore, we give motivic interpretations of them by investigating a relationship with motivic iterated integrals given by Goncharov. At this point, it is important to consider the base expansion and the Galois invariant part of the space of motivic iterated integrals. Finally, we denote that a motivic interpretation of the alternating multiple mixed values can be given by the same method. Our results also extend a part of author’s previous work.
1. Introduction
The multiple zeta values (MZVs, also called the Euler-Zagier type MZVs) are defined by
$\displaystyle\zeta(k_{1},\dots,k_{d})=\sum_{0<n_{1}<\dots<n_{d}}\frac{1}{n_{1}^{k_{1}}\dots n_{d}^{k_{d}}}$ (2)
for $k_{1},\dots,k_{d-1}\in\mathbb{Z}_{\geq{1}}$, and $k_{d}\in\mathbb{Z}_{\geq{2}}$. A pair $\mathbf{k}=(k_{1},\dots,k_{d})\in(\mathbb{Z}_{\geq{1}})^{d}$ with $k_{d}>1$ is called the (admissible) index of the weight $|\mathbf{k}|:=k_{1}+\dots+k_{d}$. It is known that there are many $\mathbb{Q}$-linear relations along MZVs ([9],[12],[13] et al.), and in [17], Zagier conjectured that the dimension of the $\mathbb{Q}$-linear space $Z_{k}$ spanned by MZVs of weight $k$ is
$\displaystyle\dim_{\mathbb{Q}}Z_{k}=d_{k},$ (3)
where
$\displaystyle\sum_{k\geq 0}d_{k}t^{k}=\frac{1}{1-t^{2}-t^{3}}.$ (4)
This conjecture has not yet been proved as of March 2024, but Goncharov ([7]), Terasoma ([15]), Deligne–Goncharov ([4]) showed
$\displaystyle\dim_{\mathbb{Q}}Z_{k}\leq d_{k}$ (5)
by using the category of the mixed Tate motives over $\mathbb{Q}$. At that time, the iterated integral representation given below was important to give a motivic interpretation of MZVs. The iterated integral representation of MZVs is said to have been first mentioned by M. Kontsevich (see [17, p.510]) and is given by
$\displaystyle\zeta(k_{1},\dots,k_{d})=\int_{0<x_{1}<\dots<x_{k}<1}\prod_{j=1}^{k}\eta_{j}(x_{j}),$ (6)
where
$\eta_{j}(x)=\begin{cases}\omega_{1}=\omega_{1}(x):=\frac{dx}{1-x},&j=k_{1}+\dots+k_{s-1}+1\text{ for some }s=1,\dots,d,\\ \omega_{0}=\omega_{0}(x):=\frac{dx}{x},&\text{otherwise}.\end{cases}$
We study the special values defined by iterated integrals. In general, for 1-forms $\eta_{1},\dots,\eta_{k}$ on a manifold $M$ and a smooth path $\gamma:[0,1]\rightarrow M$, we define
$\displaystyle\int_{\gamma}\eta_{1}\cdots\eta_{k}:=\int_{0<t_{1}<\cdots<t_{k}<1}\gamma^{*}\eta_{1}(t_{1})\cdots\gamma^{*}\eta_{k}(t_{k}),$ (7)
where $\gamma^{*}\eta_{j}\leavevmode\nobreak\ (j=1,\dots,k)$ is the pullback of $\eta_{j}$ along $\gamma$. For example, if $M=\mathbb{P}^{1}(\mathbb{C})\setminus\{0,1,\infty\}$, we have the iterated integral representations of the multiple zeta values:
$\displaystyle\zeta(k_{1},\dots,k_{r})=\int_{\mathrm{dch}}\eta_{1}\cdots\eta_{k},$ (8)
where $k=k_{1}+\cdots+k_{r}$, $\mathrm{dch}:[0,1]\rightarrow\mathbb{P}^{1}(\mathbb{C}),\mathrm{dch}(t):=t$, and
$\eta_{j}:=\begin{cases}\omega_{1},&j=k_{1}+\cdots k_{s}+1\text{ for some }s=0,\dots,r-1,\\ \omega_{0},&\text{otherwise}.\end{cases}$
In this paper, we consider iterated integrals on some specific algebraic curves and give their motivic interpretations. We also study explicit formulas for some iterated integrals. This paper is also regarded as a sequel of our previous work ([14]).
1.1. Main Results
Let $f\in\mathbb{Q}[X,Y,Z]$ be an irreducible homogeneous polynomial such that
$X_{f}:=\left\{[X:Y:Z]\in\mathbb{P}^{2}(\mathbb{C})\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ f(X,Y,Z)=0\right\}$
is a geometrically connected and projective smooth curve over $\mathbb{Q}$. We define $\varphi_{f}:X_{f}\rightarrow\mathbb{P}^{1}(\mathbb{C}),\varphi_{f}([X:Y:Z]):=[X:Z]$ and consider a local coordinates $x=\frac{X}{Z}$ and $y=\frac{Y}{Z}$ for $Z\neq 0$. Put $Y_{f}:=X_{f}\setminus\{x=0,1,\infty\}$. Let $\mathrm{dch}_{f}:[0,1]\rightarrow X_{f}$ be a smooth path which satisfies $\varphi_{f}\circ\mathrm{dch}_{f}=\mathrm{dch}$. Fix a $\mathbb{Q}$-basis $B_{f}$ of the first algebraic de Rham cohomology $H^{1}_{\mathrm{dR}}(Y_{f}/\mathbb{Q})$ (cf. [11]).
Definition 1.1.
For $\eta_{1},\dots,\eta_{k}\in B_{f}$, we define
$\displaystyle\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k}):=\int_{\mathrm{dch}_{f}}\eta_{1}\cdots\eta_{k}.$ (9)
We consider only the case when the equation 9 converges and set
$\displaystyle B_{f}^{(k)}:=\left\{(\eta_{1},\dots,\eta_{k})\in(B_{f})^{k}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})\text{ converges}\right\}.$ (10)
We call $k$ the weight of $(\eta_{1},\dots,\eta_{k})\in B_{f}^{(k)}$, and let $Z_{f}^{(k)}$ be the $\mathbb{Q}$-linear space spanned by $\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})$ for $(\eta_{1},\dots,\eta_{k})\in B_{f}^{(k)}$. When $k=0$, we interpret $\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})=1$ and set $Z_{f}^{(k)}=\mathbb{Q}$.
In this paper, we consider the following case when $f\in\{g,h\}$ where
  1. (I)
    $g(X,Y,Z):=X^{2}+Y^{2}-Z^{2};$
  2. (II)
    $h(X,Y,Z):=X^{2}+XY+Y^{2}-Z^{2}.$
The main result of this paper is to give a motivic interpretation for these values by using motivic iterated integrals. To explain the claim, we need more notation. Fix an algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$. For a number field $F(\subseteq\overline{\mathbb{Q}})$ and a $\mathbb{Q}$-vector subspace $R\subseteq F^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$, let $\mathcal{A}_{F,R}$ be the fundamental Hopf algebra of $R$ and $\mathcal{H}_{F,R}:=\mathcal{A}_{F,R}\otimes_{\mathbb{Q}}\mathbb{Q}[\tau]$. Let $\mathrm{per}:\mathcal{H}_{F,R}\rightarrow\mathbb{C}$ be the period map. Then the Galois group $\mathrm{Gal}(F/\mathbb{Q})$ acts on $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ (see Section 4 for details).
We apply them to our case for $f\in\{g,h\}$. We define a number field $F_{f}$ by
$F_{f}:=\begin{cases}\mathbb{Q}(\xi_{4}),&f=g,\\ \mathbb{Q}(\xi_{6}),&f=h,\\ \end{cases}$
where $\xi_{N}:=\exp\left(\frac{2\pi\sqrt{-1}}{N}\right)$. We introduce a $\mathbb{Q}$-vector subspace $R_{f}\subseteq F_{f}^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ which is defined by
$R_{f}:=\begin{cases}\langle 2\rangle_{\mathbb{Q}},&f=g,\\ \langle 2,3\rangle_{\mathbb{Q}},&f=h.\end{cases}$
Also, we put $G_{f}:=\mathrm{Gal}(F_{f}/\mathbb{Q})$. Then, we have the following theorem.
Theorem 1.2.
For each $f\in\{g,h\}$ and $(\eta_{1},\dots,\eta_{k})\in B_{f}^{(k)}$, there exists an element
$\mathrm{I}^{\mathfrak{m}}_{f}(\eta_{1}\cdots\eta_{k})\in\left(\mathcal{H}_{F_{f},R_{f}}^{(k)}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}$
such that $\mathrm{per}(\mathrm{I}^{\mathfrak{m}}_{f}(\eta_{1}\cdots\eta_{k}))=\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})$.
As an application of Theorem 1.2, we have an upper bounds of the dimension of $Z_{f}^{(k)}$ for each $f\in\{g,h\}$.
Corollary 1.3.
For $f\in\{g,h\}$, it holds
$\displaystyle\dim_{\mathbb{Q}}Z_{f}^{(k)}\leq D_{f}^{(k)}$ (11)
where the integer $D_{f}^{(k)}$ is defined by
$\displaystyle\sum_{k=0}^{\infty}D_{f}^{(k)}t^{k}=\begin{cases}\frac{1}{1-2t},&f=g,\\ \frac{1}{1-3t+t^{2}},&f=h.\end{cases}$ (12)
We will give the value of $D_{f}^{(k)}$ and conjectural value of $\dim_{\mathbb{Q}}$ after Corollary 5.10 obtained by numerical experiments.
Our (motivic) periods are coming from the motives defined over $\mathbb{Q}$. However, if we extend the base field $\mathbb{Q}$ suitably, they are regarded as the (motivic) periods coming from the category of the mixed Tate motives over cyclotomic fields which are well-known objects ([8]). Thus, a key, as in previous paper [14], is to consider the Galois descent of the (motivic) periods after a suitable base extension.
We organize this paper as follows. In Section 2, we set up some notations and definitions. In Section 3, we explain some basic properties of our iterated integrals and give some explicit formulas. These formulas shows our periods can be written as linear combinations of well known periods and it reads us to seek their motivic interpretations, which are the contents in Sections 4 and 5. In Section 4, we summarize the basic properties of the motivic iterated integrals given by Deligne, Glanois, and Goncharov to give motivic interpretations of the iterated integrals on our algebraic curves. In Section 5, we give motivic interpretations and investigate their algebraic structures. This interpretation explains the appearance of $\pi$, $\log 2$, $\log 3$, $L(k,\chi_{-3})$, and Riemann zeta values in the explicit formulas in Example 3.1-(3), Propositions 3.5 and 3.7. We also notice about the direct sum decomposition of the $\mathbb{Q}$-linear space $Z_{f}^{(k)}$. In Section 5.4, we describe the relationship between our periods and special values called alternating multiple mixed values, defined by Xu–Yan–Zhao [16] and further studied by Charlton [1].
We remark that we considered in [14] only the values given by iterated integrals with 33 33 elements of $B_{g}$ and we consider a more wider class in this paper.
Acknowledgement
The author would like to express his deepest gratitude to his supervisor, Professor T. Yamauchi. In particular, we follow his idea in a part of the computation in Section 3. The author would like to thank Professor H. Furusho of Nagoya University for inviting me to his monthly seminar and giving me various advice of my research. The author would also like to thank Professor M. Kaneko of Kyushu University for inviting me to the multiple zeta seminar and discussing multiple $\widetilde{T}$-values. The author is grateful to S. Charlton of the Max Planck Institute for Mathematics for discussing alternating multiple mixed values. The author would like to thank the WISE Program for AI Electrics of Tohoku University for their financial support to carry out our study.
Notation
Throughout this paper, we use the following notation unless otherwise mentioned.
Notation 1.4.
  • For a finite set $S$, let #$\#S$ be the cardinality of $S$.
  • Kronecker delta: for $r,s\in\mathbb{Z}$, put $\delta_{r,s}:=\begin{cases}1,&r=s,\\ 0,&r\neq s.\end{cases}$
  • Binomial coefficients: for $k,r\in\mathbb{Z}_{\geq{0}}$, put $\binom{k}{r}=\begin{cases}\frac{k!}{r!(k-r)!},&k\geq r,\\ 0,&k<r.\end{cases}$
  • For $n,r,s\in\mathbb{Z}_{\geq{1}}$, let $S_{n}$ be the $n$-th symmetry group, and define $(r,s)$-shuffle $S_{r,s}$ by
    $S_{r,s}:=\left\{\delta\in S_{r+s}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}\delta^{-1}(1)<\dots<\delta^{-1}(r)\\ \delta^{-1}(r+1)<\dots<\delta^{-1}(r+s)\end{array}\right\}.$
  • For $N\in\mathbb{Z}_{\geq{1}}$, put $\xi_{N}:=\exp\left(\frac{2\pi\sqrt{-1}}{N}\right)$.
  • For $N\in\mathbb{Z}_{\geq{1}}$, put $\mu_{N}:=\left\{\xi_{N}^{r}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ r=0,1,\dots,N-1\right\}$ and $\widetilde{\mu}_{N}:=\mu_{N}\cup\{0\}$.
  • Path reversal: for a piecewise smooth path $\gamma:[0,1]\rightarrow\mathbb{C}$, let $\gamma^{-1}:[0,1]\rightarrow\mathbb{C}$ be
    $\gamma^{-1}(t):=\gamma(1-t).$
  • Path connection: for piecewise smooth paths $\gamma_{1},\gamma_{2}:[0,1]\rightarrow\mathbb{C}$ satisfying $\gamma_{1}(1)=\gamma_{2}(0)$, let $\gamma_{1}\gamma_{2}:[0,1]\rightarrow\mathbb{C}$ be
    $(\gamma_{1}\gamma_{2})(t):=\begin{cases}\gamma_{1}(2t),&0<t<1/2,\\ \gamma_{2}(2t-1),&1/2<t<1.\end{cases}$
  • For $p,q\in\mathbb{C}$, we define the straight path $\mathrm{dch}_{p,q}$ on $\mathbb{C}$ by
    $\mathrm{dch}_{p,q}:[0,1]\rightarrow\mathbb{C},\leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{dch}_{p,q}(t)=p+t(q-p).$
    In particular, for $p=0$ and $q=1$, we put $\mathrm{dch}:=\mathrm{dch}_{0,1}$.
  • For $p,q_{1},q_{2}\in\mathbb{C}$ with $|p-q_{1}|=|p-q_{2}|=r$, we define the arc $C_{p}(q_{1},q_{2})$ on $\mathbb{C}$ by
    $C_{p}(q_{1},q_{2}):[0,1]\rightarrow\mathbb{C},\leavevmode\nobreak\ \leavevmode\nobreak\ C_{p}(q_{1},q_{2})(t)=p+re^{\sqrt{-1}(\arg(q_{1})+t(\arg(q_{2}/q_{1})))}.$
  • For a number field $F$, we put $r_{1}(F):=\#\{\text{real places}\}$ and $r_{2}(F):=\#\{\text{complex places}\}$.
  • For an $F$-linear space $V$, let $T(V)=\bigoplus_{k=0}^{\infty}V^{\otimes k}$ be the tensor algebra of $V$.
  • For a field $F$ and a set $S$, let $F\langle S\rangle:=F\langle e_{s}\leavevmode\nobreak\ |\leavevmode\nobreak\ s\in S\rangle$ be the non-commutative $\mathbb{Q}$-algebra generated by the formal basis $e_{s}\leavevmode\nobreak\ (s\in S)$.
2. Preparations
First, we prepare some notation for iterated integrals on $\mathbb{C}$.
Definition 2.1.
Let $\gamma:[0,1]\rightarrow\mathbb{C}$ be a piecewise smooth path on $\mathbb{C}$ and $a_{1},\dots,a_{k}\in\mathbb{C}\setminus\gamma((0,1))$ with $p:=\gamma(0)\neq a_{1},q:=\gamma(1)\neq a_{k}$. Then, we define
$\displaystyle\mathrm{I}_{\gamma}(p;a_{1},\dots,a_{k};q):=\int_{\gamma}\frac{\leavevmode\nobreak\ dx}{x-a_{1}}\dots\frac{\leavevmode\nobreak\ dx}{x-a_{k}}.$ (13)
For the moment, we do not ask if the above iterated integral converges.
In addition, for each element $w=e_{a_{1}}\dots e_{a_{k}}\in\mathbb{Q}\langle\mathbb{C}\rangle$ with $a_{1}\neq p,a_{k}\neq q$, we define
$\displaystyle\mathrm{I}_{\gamma}(p;w;q):=\mathrm{I}_{\gamma}(p,a_{1},\dots,a_{k};q)$ (14)
and extend $\mathrm{I}_{\gamma}(p;w;q)$$\mathbb{Q}$-linearly to any $w\in\mathbb{Q}\langle\mathbb{C}\rangle$.
Remark 2.2.
The right hand side of the equation 13 diverges if $p=a_{1}$, $q=a_{k}$, or $a_{1},\dots,a_{k}\in\gamma((0,1))$. However, we can regularize $\mathrm{I}_{\gamma}(p;a_{1},\dots,a_{k};q)\in\mathbb{C}$ in these cases as in [14, p.6, Definition 2.0.6]. For example, if $k=1$, then it holds
$\displaystyle\exp(\mathrm{I}_{\mathrm{dch}_{a,c}}(a;b;c))=\widetilde{\mathrm{I}}(a,b,c),$ (15)
where
$\displaystyle\widetilde{\mathrm{I}}(a,b,c):=\begin{cases}\frac{c-b}{a-b},&a\neq b\text{ and }b\neq c,\\ c-b,&a=b\text{ and }b\neq c,\\ (a-b)^{-1},&a\neq b\text{ and }b=c,\\ 1&a=b=c.\end{cases}$ (16)
Definition 2.3.
For $k_{1},\dots,k_{d}\in\mathbb{Z}_{\geq{1}}$ and $\alpha_{1},\dots,\alpha_{d}\in\mu_{N}$ with $(k_{d},\alpha_{d})\neq(1,1)$,
$\displaystyle L\left(\begin{matrix}k_{1},\dots,k_{d}\\ \alpha_{1},\dots,\alpha_{d}\end{matrix}\right):=(-1)^{d}\mathrm{I}_{\mathrm{dch}}(0;a_{1}^{-1},\{0\}^{k_{1}-1},\dots,\alpha_{d}^{-1},\{0\}^{k_{d}-1};1)$ (19)
is called the (shuffle type) multiple $L$-value of level $N$.
Next, we consider iterated integrals in Definition 1.1 for our algebraic curves.
Example 2.4.
If $f(X,Y,Z)=X^{2}+Y^{2}-Z^{2}$, we can take a $\mathbb{Q}$-basis as $B_{f}=\left\{\frac{\leavevmode\nobreak\ dx}{x},\frac{\leavevmode\nobreak\ dx}{1-x},\frac{\leavevmode\nobreak\ dx}{y},\frac{\leavevmode\nobreak\ dx}{xy}\right\}$. Then, if we restrict the 11 11-forms to the subset $\left\{\frac{\leavevmode\nobreak\ dx}{x},\frac{\leavevmode\nobreak\ dx}{1-x},\frac{\leavevmode\nobreak\ dx}{y}\right\}\subseteq B_{f}$, the iterated integrals equation 9 give $F_{2}$-MZVs considered in [14].
Now, we consider the cases of $g(X,Y,Z)=X^{2}+Y^{2}-Z^{2}$ and $h(X,Y,Z)=X^{2}+XY+Y^{2}-Z^{2}$. Note that for each $f\in\{g,h\}$, the geometric genus of $X_{f}$ is 00 0. Further, $X_{g}$ is smooth over $\mathbb{Z}[1/2]$ and $X_{h}$ is smooth over $\mathbb{Z}[1/3]$.
An explicit $\mathbb{Q}$-basis of $H_{\mathrm{dR}}^{1}(Y_{f}/\mathbb{Q})$. For each $X_{f}$ ($f\in\{g,h\}$), we specify $B_{f}$ as follows.
  1. (I)
    $g(X,Y,Z)=X^{2}+Y^{2}-Z^{2}$: There is an isomorphism
    $\displaystyle\phi_{g}:\mathbb{P}^{1}(\mathbb{C})\overset{\sim}{\longrightarrow}X_{g},\phi_{g}(\lambda):=(x,y)=\left(\frac{2\lambda}{1+\lambda^{2}},\frac{1-\lambda^{2}}{1+\lambda^{2}}\right).$ (20)
    The points on $X_{g}$ with $x=0,1,\infty$ correspond to the following 55 55 points: $(x,y)=\left(0,\pm 1\right),(1,0)$ and $[X:Y:Z]=\left[1:\pm\sqrt{-1}:0\right]$. The dimension of $H^{1}_{\mathrm{dR}}(Y_{g}/\mathbb{Q})$ is 44 44 and we can take $B_{g}$ as
    $B_{g}=\left\{\omega_{0}:=\frac{\leavevmode\nobreak\ dx}{x},\omega_{1}:=\frac{\leavevmode\nobreak\ dx}{1-x},\omega_{2}:=\frac{\leavevmode\nobreak\ dx}{y},\omega_{3}:=\frac{\leavevmode\nobreak\ dx}{xy}\right\}.$
    The set $B_{g}^{(k)}$ defined by the equation 10 becomes
    $B_{g}^{(k)}=\{(\eta_{1},\dots,\eta_{k})\in(B_{g})^{k}\leavevmode\nobreak\ |\leavevmode\nobreak\ \eta_{1}\neq\omega_{0},\omega_{3}\text{ and }\eta_{k}\neq\omega_{1}\}.$
  2. (II)
    $h(X,Y,Z)=X^{2}+XY+Y^{2}-Z^{2}$: There is an isomorphism
    $\displaystyle\phi_{h}:\mathbb{P}^{1}(\mathbb{C})\overset{\sim}{\longrightarrow}X_{h},\phi_{h}(\lambda):=(x,y)=\left(\frac{\lambda(\lambda+2)}{1+\lambda+\lambda^{2}},\frac{1-\lambda^{2}}{1+\lambda+\lambda^{2}}\right).$ (21)
    The points on $X_{h}$ with $x=0,1,\infty$ correspond to the following 66 66 points: $(x,y)=(0,\pm 1),(1,0),(1,-1)$ and $[X:Y:Z]=[1:\xi_{3}^{\pm 1}:0]$. The dimension of $H^{1}_{\mathrm{dR}}(Y_{h}/\mathbb{Q})$ is 55 55 and we can take $B_{h}$ as
    $B_{h}=\left\{\omega_{0}:=\frac{\leavevmode\nobreak\ dx}{x},\omega_{1}:=\frac{\leavevmode\nobreak\ dx}{1-x},\omega_{4}:=\frac{\leavevmode\nobreak\ dx}{x+2y},\omega_{5}:=\frac{\leavevmode\nobreak\ dx}{x(x+2y)},\omega_{6}:=\frac{\leavevmode\nobreak\ dx}{(1-x)(x+2y)}\right\}.$
    The set $B_{h}^{(k)}$ defined by the equation 10 becomes
    $B_{h}^{(k)}=\{(\eta_{1},\dots,\eta_{k})\in(B_{h})^{k}\leavevmode\nobreak\ |\leavevmode\nobreak\ \eta_{1}\neq\omega_{0},\omega_{5}\text{ and }\eta_{k}\neq\omega_{1},\omega_{6}\}.$
In the following, for each $f\in\{g,h\}$, we consider only the case for $(\eta_{1},\dots,\eta_{k})\in B_{f}^{(k)}$.
3. Basic properties
Now, we explain some basic properties of our target.
3.1. Relationship with known special values
First, we describe the relationship with known special values.
Example 3.1.
Take $f\in\{g,h\}$ and $k,k_{1},\dots,k_{d}\in\mathbb{Z}_{\geq{1}}$ with $k_{d}\geq 2$.
  1. (1)
    By iterated integral representation of multiple zeta values (equation 8), it holds
    $\mathrm{I}_{f}(\omega_{1}\omega_{0}^{k_{1}-1}\cdots\omega_{1}\omega_{0}^{k_{d}-1})=\zeta(k_{1},\dots,k_{d}).$
  2. (2)
    By substituting the integral variable $x=\frac{2\lambda}{1+\lambda^{2}}$, we have
    $\displaystyle\mathrm{I}_{g}(\omega_{2}\omega_{3}^{k_{1}-1}\cdots\omega_{2}\omega_{3}^{k_{d}-1})=\int_{0}^{1}\frac{2\leavevmode\nobreak\ d\lambda}{1+\lambda^{2}}\left(\frac{\leavevmode\nobreak\ d\lambda}{\lambda}\right)^{k_{1}-1}\cdots\frac{2\leavevmode\nobreak\ d\lambda}{1+\lambda^{2}}\left(\frac{\leavevmode\nobreak\ d\lambda}{\lambda}\right)^{k_{d}-1}.$ (22)
    This matches the iterated integral representation of multiple $\widetilde{T}$-values
    $\displaystyle\widetilde{T}(k_{1},\dots,k_{n}):=2^{n}\sum_{\begin{subarray}{c}0<m_{1}<\cdots<m_{n}\\ m_{j}\equiv j\text{ mod }2\end{subarray}}\frac{(-1)^{(m_{n}-n)/2}}{m_{1}^{k_{1}}\cdots m_{n}^{k_{n}}},$ (25)
    which is defined by Kaneko–Tsumura [10, p.4, Proposition 2.1]. Then, it holds
    $\displaystyle\mathrm{I}_{g}(\omega_{2}\omega_{3}^{k_{1}-1}\cdots\omega_{2}\omega_{3}^{k_{d}-1})=\widetilde{T}(k_{1},\dots,k_{n}).$ (26)
  3. (3)
    In our previous work ([14, p.9, Proposition 3.2.1]), we showed
    $\displaystyle\mathrm{I}_{g}(\omega_{2}\omega_{0}^{k-1})\in\mathrm{Span}_{\mathbb{Q}}\left\{\pi^{l}\cdot(\log 2)^{l_{1}}\cdot\prod_{j\geq 3\text{: odd}}\zeta(j)^{l_{j}}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}l\in\mathbb{Z}_{\geq{1}}\text{: odd},\leavevmode\nobreak\ l_{j}\in\mathbb{Z}_{\geq{0}}\leavevmode\nobreak\ (j\geq 1\text{: odd})\\ l+\sum_{j\geq 1\text{: odd}}l_{j}=k\end{array}\right\}.$ (29)
3.2. Explicit formula
Here, we summarize formulas given by the definition and the basic properties of iterated integrals.
Proposition 3.2 ([2, p.220, p233]).
For a smooth manifold $M$, 11 11-forms $\eta_{1},\dots,\eta_{k},\eta_{k+1},\dots,\eta_{k+l}$ on $M$, and piecewise smooth path $\gamma,\gamma_{1},\gamma_{2}:[0,1]\rightarrow M$ with $\gamma_{1}(1)=\gamma_{2}(0)$, the following properties hold.
  1. (1)
    (path connection)
    $\int_{\gamma_{1}\gamma_{2}}\eta_{1}\cdots\eta_{k}=\sum_{j=0}^{k}\int_{\gamma_{1}}\eta_{1}\cdots\eta_{j}\int_{\gamma_{2}}\eta_{j+1}\cdots\eta_{k}.$
  2. (2)
    (path reversal)
    $\int_{\gamma^{-1}}\eta_{1}\cdots\eta_{k}=(-1)^{k}\int_{\gamma}\eta_{k}\cdots\eta_{1}.$
  3. (3)
    (shuffle relation)
    $\int_{\gamma}\eta_{1}\cdots\eta_{k}\int_{\gamma}\eta_{k+1}\cdots\eta_{k+l}=\sum_{\sigma\in S_{k,l}}\int_{\gamma}\eta_{\sigma(1)}\cdots\eta_{\sigma(k+l)}.$
  4. (4)
    (homotopy invariance) If each of $\eta_{1},\dots,\eta_{k}$ is closed and satisfy $\eta_{j}\wedge\eta_{j+1}=0\leavevmode\nobreak\ (j=1,\dots,k-1)$, the value $\int_{\gamma}\eta_{1}\cdots\eta_{k}$ only depends on the homotopy class of $\gamma$.
By Proposition 3.2-(3), we have the following proposition.
Proposition 3.3.
For each $f\in\{g,h\}$ and $(\varphi_{1},\dots,\varphi_{k})\in B_{f}^{(k)},(\varphi_{k+1},\dots,\varphi_{k+l})\in B_{f}^{(l)}$, it holds
$\displaystyle\mathrm{I}_{f}(\varphi_{1}\cdots\varphi_{k})\mathrm{I}_{f}(\varphi_{k+1}\cdots\varphi_{k+l})=\sum_{\sigma\in S_{k,l}}\mathrm{I}_{f}(\varphi_{\sigma(1)}\cdots\varphi_{\sigma(k+l)}).$ (30)
In particular, it holds that
  1. (1)
    for each $f\in\{g,h\}$, $k\in\mathbb{Z}_{\geq{0}}$, and $(\varphi)\in B_{f}^{(1)}$,
    $\displaystyle\mathrm{I}_{f}(\varphi^{k})=\frac{1}{k!}\mathrm{I}_{f}(\varphi)^{k};$ (31)
  2. (2)
    for each $f\in\{g,h\}$, $k\in\mathbb{Z}_{\geq{0}}$, $j=0,\dots,k-2$, and $\varphi,\eta\in B_{f}$ with ($(\{\varphi\}^{j},\eta,\{\varphi\}^{k-j-1})\in B_{f}^{(k)}$,
    $\displaystyle\mathrm{I}_{f}(\varphi^{j}\eta\varphi^{k-j-1})=\frac{1}{k-j-1}\left\{\mathrm{I}_{f}(\varphi^{j}\eta\varphi^{k-j-2})\mathrm{I}_{f}(\varphi)-(j+1)\mathrm{I}_{f}(\varphi^{j+1}\eta\varphi^{k-j-2})\right\}.$ (32)
Now, we give the iterated sum representation of our target.
Proposition 3.4.
For each $f\in\{g,h\}$ and $(\varphi_{1},\dots,\varphi_{k})\in B_{f}^{(k)}$, it holds
$\displaystyle\mathrm{I}_{f}(\varphi_{1}\cdots\varphi_{k})=\sum_{0<n_{1}\leq\dots\leq n_{k}}\prod_{j=1}^{k}\frac{c(n_{j}-n_{j-1},\varphi_{j})}{n_{j}},$ (33)
where $n_{0}=0$ and
$\displaystyle c(n,\omega_{0})$ $\displaystyle=\delta_{n,0},$ $\displaystyle c(n,\omega_{1})$ $\displaystyle=1-\delta_{n,0}$ (34)
$\displaystyle c(n,\omega_{2})$ =$\displaystyle=\begin{cases}\frac{1}{2^{n-1}}\binom{n-1}{\frac{n-1}{2}},&n\text{ is odd},\\ 0,&n\text{ is even},\end{cases}$ $\displaystyle c(n,\omega_{3})$ =$\displaystyle=\begin{cases}\frac{1}{2^{n}}\binom{n}{\frac{n}{2}},&n\text{ is even},\\ 0,&n\text{ is odd},\end{cases}$ (35)
$\displaystyle c(n,\omega_{4})$ =$\displaystyle=\begin{cases}\left(\frac{\sqrt{3}}{4}\right)^{n-1}\binom{n-1}{\frac{n-1}{2}},&n\text{ is odd},\\ 0,&n\text{ is even},\end{cases}$ $\displaystyle c(n,\omega_{5})$ =$\displaystyle=\begin{cases}\left(\frac{\sqrt{3}}{4}\right)^{n}\binom{n}{\frac{n}{2}},&n\text{ is even},\\ 0,&n\text{ is odd},\end{cases}$ (36)
$\displaystyle c(n,\omega_{6})$ $\displaystyle=\frac{1}{2}\sum_{m=0}^{\lfloor\frac{n-1}{2}\rfloor}\left(\frac{3}{16}\right)^{m}\binom{2m}{m}$ (37)
for $n\in\mathbb{Z}_{\geq{0}}$.
Proof ..
It is known that
$\displaystyle\int_{\mathrm{dch}}\varphi_{1}\cdots\varphi_{k}=\sum_{0<n_{1}\leq\dots\leq n_{k}}\prod_{j=1}^{k}\frac{c(n_{j}-n_{j-1},\varphi_{j})}{n_{j}},$ (38)
where $\varphi_{j}=\sum_{n=0}^{\infty}c(n,\varphi_{j})x^{n-1}\leavevmode\nobreak\ dx$ (see [14, p.8, Proposition 3.1.3]). It follows equation 33 since $\omega_{s}=\sum_{n=0}^{\infty}c(n,\omega_{s})x^{n-1}\leavevmode\nobreak\ dx$ for each $s=0,1,\dots,6$. ∎
Now, we investigate a special case of which our target can be written in terms of well-known special values as logarithms, Riemann zeta values, and Dirichlet $L$-values.
Proposition 3.5.
For each positive integer $m\geq 2$, it holds
$\displaystyle\mathrm{I}_{g}(\omega_{2}^{2m-1}\omega_{0})$ $\displaystyle=-2^{1-2m}\sum_{l=1}^{m-1}\frac{(-1)^{l}}{(2m-2l-1)!}(2^{-2l}-1)\zeta(2l+1)\pi^{2m-2l-1}$ (39)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{(2m-1)!}\log 2\left(\frac{\pi}{2}\right)^{2m-1}.$ (40)
Also, for each positive integer $m\geq 1$, it holds
$\displaystyle\mathrm{I}_{g}(\omega_{2}^{2m}\omega_{0})$ $\displaystyle=-2^{-2m}\sum_{l=1}^{m-1}\frac{(-1)^{l}}{(2m-2l)!}(2^{-2l}-1)\zeta(2l+1)\pi^{2m-2l}$ (41)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +(-1)^{m}2^{-2m}(2-2^{-2m})\zeta(2m+1)+\frac{1}{(2m)!}\log 2\left(\frac{\pi}{2}\right)^{2m}.$ (42)
In particular,
$\displaystyle\mathrm{I}_{g}(\omega_{2}^{k-1}\omega_{0})\in\mathrm{Span}_{\mathbb{Q}}\left\{\log 2\cdot\pi^{k-1},\zeta(2l+1)\pi^{k-2l-1}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ l=1,\dots,\left\lfloor\frac{k-1}{2}\right\rfloor\right\}.$ (43)
Also, for $k\geq 2$ and $2\displaystyle 1\leq j\leq k-$ 2, it holds
$\displaystyle\mathrm{I}_{g}(\omega_{2}^{j}\omega_{0}\omega_{2}^{k-j-1})\in\mathrm{Span}_{\mathbb{Q}}\left\{\zeta(2l+1)\pi^{k-2l-1}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ l=1,\dots,\left\lfloor\frac{k-1}{2}\right\rfloor\right\}.$ (44)
To prove this explicit formula, we read the following lemma.
Lemma 3.6.
If $z_{0}\in\mathbb{C}$ satisfies $|z_{0}|=1$ and $-\frac{\pi}{2}\leq\arg(z_{0})\leq\frac{\pi}{2}$, they hold
[…]
where $\gamma=C_{0}(1,z_{0})$ and
$\displaystyle\mathrm{Li}_{k}(z_{0}):=-\mathrm{I}_{\mathrm{dch}_{0,z}}(0;1,\{0\}^{k-1};z_{0})=\sum_{n=1}^{\infty}\frac{z_{0}^{n}}{n^{k}}.$ (47)
Proof ..
First we show the equation 45. We have
[…]
where $\gamma_{1}=\mathrm{dch}_{0,1-\varepsilon}$, $\gamma_{2}=C_{1}(1-\varepsilon,1+\varepsilon\sqrt{-1})$, and $\gamma_{3}=C_{0}(1+\varepsilon\sqrt{-1},z_{2})$ for sufficiently small $\varepsilon>0$ (see FIGURE 1).
[…]
Рис. 1. integral paths
By $\varepsilon\searrow 0$ and calculating the regularization of the iterated integral (Remark 2.2), they hold
$\displaystyle\int_{\gamma_{1}}\frac{\leavevmode\nobreak\ dz_{1}}{1-z_{1}}=0$ (50)
$\displaystyle\int_{\gamma_{2}}\frac{\leavevmode\nobreak\ dz_{1}}{1-z_{1}}=-\frac{\pi\sqrt{-1}}{2}.$ (51)
Also,
[…]
Then, we have the equation 45. Next, we show the equation 46. We have
[…]
where $\gamma_{3}=\mathrm{dch}$ and $\gamma_{4}=C_{0}(1,z_{2})$. Here,
$\displaystyle\int_{\gamma}\frac{\leavevmode\nobreak\ dz_{1}}{1+z_{1}}=\log 2.$ (55)
On the other hands, it holds
[…]
$\displaystyle=-\int_{\mathrm{dch}_{-1,-z_{0}}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}$ (57)
$\displaystyle=-\int_{\gamma_{5}\gamma_{6}\gamma_{7}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z},$ (58)
where $\gamma_{5}=\mathrm{dch}_{-1,-\varepsilon}$, $\gamma_{6}=C_{0}(-\varepsilon,\varepsilon\arg(-z_{0}))$, and $\gamma_{7}=\mathrm{dch}_{\varepsilon\arg(-z_{0}),-z_{0}}$ for sufficiently small $\varepsilon>0$ (see FIGURE 2).
[…]
Рис. 2. integral paths
Then, by $\varepsilon\searrow 0$, Proposition 3.2 (1), and
$\displaystyle\int_{\gamma_{7}}\frac{\leavevmode\nobreak\ dz}{z}=\int_{\gamma_{6}}\frac{\leavevmode\nobreak\ dz}{1-z}=0,$ (59)
we have
$\displaystyle\int_{\gamma_{5}\gamma_{6}\gamma_{7}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}=\int_{\gamma_{5}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}+\int_{\gamma_{5}}\frac{\leavevmode\nobreak\ dz}{1-z}\int_{\gamma_{6}}\frac{\leavevmode\nobreak\ dz}{z}+\int_{\gamma_{7}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}.$ (60)
Here, each terms can be calculated as
$\displaystyle\int_{\gamma_{5}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}$ $\displaystyle=\int_{\gamma_{5}^{-1}}\frac{\leavevmode\nobreak\ dz}{z}\frac{\leavevmode\nobreak\ dz}{1-z}$ (61)
$\displaystyle=\int_{\mathrm{dch}_{0,-1}}\frac{\leavevmode\nobreak\ dz}{z}\int_{\mathrm{dch}_{0,-1}}\frac{\leavevmode\nobreak\ dz}{1-z}-\int_{\mathrm{dch}_{0,-1}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}$ (62)
$\displaystyle=-\mathrm{Li}_{2}(-1),$ (63)
$\displaystyle\int_{\gamma_{5}}\frac{\leavevmode\nobreak\ dz}{1-z}\int_{\gamma_{6}}$ $\displaystyle\frac{\leavevmode\nobreak\ dz}{z}=\log 2\log z_{0},$ (64)
$\displaystyle\int_{\gamma_{7}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}$ $\displaystyle=\int_{\mathrm{dch}_{0,-z_{0}}}\frac{\leavevmode\nobreak\ dz}{1-z}\frac{\leavevmode\nobreak\ dz}{z}=\mathrm{Li}_{2}(-z_{0}),$ (65)
then we have the equation 46. ∎
Proof of Proposition 3.5..
By the definition and the substitution $x=\sin t$, for $k\geq 2$,
$\displaystyle\mathrm{I}_{g}(\omega_{2}^{k-1}\omega_{0})$ $\displaystyle=\int_{0}^{1}\underbrace{\frac{\leavevmode\nobreak\ dx}{\sqrt{1-x^{2}}}\cdots\frac{\leavevmode\nobreak\ dx}{\sqrt{1-x^{2}}}}_{k-1}\frac{\leavevmode\nobreak\ dx}{x}$ (66)
$\displaystyle=\int_{0}^{\frac{\pi}{2}}\underbrace{\leavevmode\nobreak\ dt\cdots\leavevmode\nobreak\ dt}_{k-1}\cot t\leavevmode\nobreak\ dt$ (67)
$\displaystyle=\frac{1}{(k-1)!}\int_{0}^{\frac{\pi}{2}}t^{k-1}\cot t\leavevmode\nobreak\ dt$ (68)
$\displaystyle=\frac{1}{(k-1)!}\left\{\left[t^{k-1}\log\sin t\right]_{0}^{\frac{\pi}{2}}-(k-1)\int_{0}^{\frac{\pi}{2}}t^{k-2}\log\sin t\leavevmode\nobreak\ dt\right\}$ (69)
$\displaystyle=-\frac{1}{(k-2)!}\int_{0}^{\frac{\pi}{2}}t^{k-2}\log\sin t\leavevmode\nobreak\ dt.$ (70)
Then, if we put $z=e^{\sqrt{-1}t}$,
$\displaystyle\int_{0}^{\frac{\pi}{2}}t^{k-2}\log\sin t\leavevmode\nobreak\ dt=\frac{1}{\sqrt{-1}^{k-1}}\int_{\gamma}\log^{k-2}z\log\frac{1-z^{2}}{-2\sqrt{-1}z}\frac{\leavevmode\nobreak\ dz}{z},$ (71)
where $z=C_{0}(1,\sqrt{-1})$. Since $0\leq\arg(z)\leq\frac{\pi}{2}$ for any $z\in\gamma([0,1])$, we have
$\displaystyle-\frac{\pi}{2}\leq\arg(1-z)\leq-\frac{\pi}{4},$ $\displaystyle 0\leq\arg(1+z)\leq\frac{\pi}{4},$ $\displaystyle-\frac{\pi}{2}\leq\arg(-2\sqrt{-1}z)\leq 0.$ (72)
Then, by Lemma 3.6 and the partial integration, we have
$\displaystyle\int_{\gamma}\log^{k-2}z\log\frac{1-z^{2}}{-2\sqrt{-1}z}\frac{\leavevmode\nobreak\ dz}{z}$ (73)
$\displaystyle=\left[\log^{k-2}z(-\mathrm{Li}_{2}(z)+\frac{\pi\sqrt{-1}}{2}\log z-\mathrm{Li}_{2}(-z)+\log 2\log z+\frac{1}{2}\log^{2}z)\right]_{1}^{\sqrt{-1}}$ (74)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -(k-2)\int_{1}^{\sqrt{-1}}\log^{k-3}z(-\mathrm{Li}_{2}(z)+\frac{\pi\sqrt{-1}}{2}\log z-\mathrm{Li}_{2}(-z)+\log 2\log z+\frac{1}{2}\log^{2}z)\frac{\leavevmode\nobreak\ dz}{z}.$ (75)
Furthermore, by repeating the partial integration, we have
$\displaystyle\int_{\gamma}\log^{k-2}z\log\frac{1-z^{2}}{-2\sqrt{-1}z}\frac{\leavevmode\nobreak\ dz}{z}$ (76)
$\displaystyle=\sum_{j=2}^{k}(-1)^{j-1}\frac{(k-2)!}{(k-j)!}\left(\frac{\pi\sqrt{-1}}{2}\right)^{k-j}\left(\mathrm{Li}_{j}(\sqrt{-1})+\mathrm{Li}_{j}(-\sqrt{-1})\right)+(-1)^{j}\frac{(k-2)!}{(k-j)!(j-1)!}\left(\frac{\pi\sqrt{-1}}{2}\right)^{k}$ (77)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +(-1)^{k}(k-2)!\left(\mathrm{Li}_{k}(1)+\mathrm{Li}_{k}(-1)\right)-\frac{\log 2}{k-1}\left(\frac{\pi\sqrt{-1}}{2}\right)^{k-1}+\frac{1}{k-1}\left(\frac{\pi\sqrt{-1}}{2}\right)^{k}-\frac{1}{k}\left(\frac{\pi\sqrt{-1}}{2}\right)^{k}.$ (78)
Since $\mathrm{I}_{g}(\omega_{2}^{k-1}\omega_{0})$ is a real number, we consider only the real part of the equation 70. Also, it is known that
$\displaystyle\mathrm{Li}_{j}(\sqrt{-1})+\mathrm{Li}_{j}(-\sqrt{-1})=2^{1-j}(2^{1-j}-1)\zeta(j);$ (79)
$\displaystyle\mathrm{Li}_{j}(1)+\mathrm{Li}_{j}(-1)=2^{1-j}\zeta(j)$ (80)
then we have
$\displaystyle\mathrm{I}_{g}(\omega_{2}^{k-1}\omega_{0})$ $\displaystyle=-\sum_{\begin{subarray}{c}j=2\\ j:\text{odd}\end{subarray}}^{k}(-1)^{\frac{1-j}{2}}\frac{1}{(k-j)!}\left(\frac{\pi}{2}\right)^{k-j}2^{1-j}(2^{1-j}-1)\zeta(j)$ (83)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1-(-1)^{k}}{2}(-1)^{\frac{1-k}{2}}2^{1-k}\zeta(k)+\frac{1}{(k-1)!}\log 2\left(\frac{\pi}{2}\right)^{k-1}.$ (84)
The equations 40 and 42 are shown by setting $k=2m$ and $k=2m+1$. Furthermore, by Proposition 3.3,
$\displaystyle\mathrm{I}_{g}(\omega_{2}^{k-2}\omega_{0}\omega_{2})=\mathrm{I}_{g}(\omega_{2}^{k-2}\omega_{0})\mathrm{I}_{g}(\omega_{2})-(k-1)\mathrm{I}_{g}(\omega_{2}^{k-1}\omega_{0}).$ (85)
Thus, by the equation 84 and $\mathrm{I}_{g}(\omega_{2})=\frac{\pi\sqrt{-1}}{2}$, the terms containing $\log 2$ vanish and we have
$\displaystyle\mathrm{I}_{g}(\omega_{2}^{k-2}\omega_{0}\omega_{2})\in\mathrm{Span}_{\mathbb{Q}}\left\{\zeta(2l+1)\pi^{k-2l-1}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ l=1,\dots,\left\lfloor\frac{k-1}{2}\right\rfloor\right\}.$ (86)
Also, by repeating the equation
$\displaystyle\mathrm{I}_{f}(\varphi^{j}\eta\varphi^{k-j-1})=\frac{1}{k-j-1}\left\{\mathrm{I}_{f}(\varphi^{j}\eta\varphi^{k-j-2})\mathrm{I}_{f}(\varphi)-(j+1)\mathrm{I}_{f}(\varphi^{j+1}\eta\varphi^{k-j-2})\right\},$ (87)
we have
$\displaystyle\mathrm{I}_{g}(\omega_{2}^{j}\omega_{0}\omega_{2}^{k-j-1})\in\mathrm{Span}_{\mathbb{Q}}\left\{\zeta(2l+1)\pi^{k-2l-1}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ l=1,\dots,\left\lfloor\frac{k-1}{2}\right\rfloor\right\}$ (88)
if $\leq j\leq k-2$. ∎
Proposition 3.7.
For each positive integer $m\geq 2$, it holds
$\displaystyle\mathrm{I}_{h}(\omega_{4}^{2m-1}\omega_{0})$ $\displaystyle=\frac{1}{3^{2m-1}}\frac{1}{(2m-1)!}\left(\frac{\pi}{\sqrt{3}}\right)^{2m-1}\frac{1}{2}\log 3$ (89)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{3^{2m-1}}\sum_{l=1}^{m}\frac{(-1)^{1-l}}{(2m-2l)!}\left(\frac{\pi}{\sqrt{3}}\right)^{2m-2l}3^{l}2^{-2l}L(2l,\chi_{-3})$ (90)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\frac{1}{3^{2m-1}}\sum_{l=1}^{m-1}\frac{(-1)^{-l}}{(2m-2l-1)!}\left(\frac{\pi}{\sqrt{3}}\right)^{2m-2l-1}3^{l}2^{-2l-1}(3^{-2l}-1)\zeta(2l+1).$ (91)
Here, $\chi_{-3}:\mathbb{Z}\rightarrow\mathbb{C}$ is the Dirichlet character defined by
$\chi_{-3}(n)=\begin{cases}1,&n\equiv 1\text{ mod $\displaystyle 3$},\\ -1&n\equiv-1\text{ mod $\displaystyle 3$},\\ 0,&n\equiv 0\text{ mod $\displaystyle 3$},\end{cases}$
and
$L(j,\chi_{-3}):=\sum_{n\geq 1}\frac{\chi_{-3}(n)}{n^{j}}$
is the Dirichlet $L$-value of $\chi_{-3}$. Also, for each positive integer $m\geq 1$, it holds
$\displaystyle\mathrm{I}_{h}(\omega_{4}^{2m}\omega_{0})$ $\displaystyle=\frac{1}{3^{2m}}\frac{1}{(2m)!}\left(\frac{\pi}{3}\right)^{2m}\frac{1}{2}\log 3+(-1)^{-m}\frac{1}{12^{m}}\zeta(2m+1)$ (92)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{3^{2m}}\sum_{l=1}^{m}\frac{(-1)^{1-l}}{(2m-2l+1)!}\left(\frac{\pi}{\sqrt{3}}\right)^{2m-2l+1}3^{l}2^{-2l}L(2l,\chi_{-3})$ (93)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\frac{1}{3^{2m}}\sum_{l=1}^{m}\frac{(-1)^{-l}}{(2m-2l)!}\left(\frac{\pi}{\sqrt{3}}\right)^{2m-2l}3^{l}2^{-2l-1}(3^{-2l}-1)\zeta(2l+1).$ (94)
In particular,
$\displaystyle\mathrm{I}_{h}(\omega_{4}^{k-1}\omega_{0})$ (95)
$\displaystyle\in\mathrm{Span}_{\mathbb{Q}}\left\{\log 3\cdot\left(\frac{\pi}{\sqrt{3}}\right)^{k-1},\zeta(2l+1)\left(\frac{\pi}{\sqrt{3}}\right)^{k-2l-1},L(2l^{\prime},\chi_{-3})\left(\frac{\pi}{\sqrt{3}}\right)^{k-2l^{\prime}}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{l}l=1,\dots,\left\lfloor\frac{k-1}{2}\right\rfloor,\\ l^{\prime}=1,\dots,\left\lfloor\frac{k}{2}\right\rfloor\end{array}\right\}.$ (98)
Also, for $k\geq 2$ and $2\displaystyle 1\leq j\leq k-$ 2, it holds
$\displaystyle\mathrm{I}_{h}(\omega_{4}^{j}\omega_{0}\omega_{4}^{k-j-1})\in\mathrm{Span}_{\mathbb{Q}}\left\{\zeta(2l+1)\left(\frac{\pi}{\sqrt{3}}\right)^{k-2l-1},L(2l^{\prime},\chi_{-3})\left(\frac{\pi}{\sqrt{3}}\right)^{k-2l^{\prime}}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{l}l=1,\dots,\left\lfloor\frac{k-1}{2}\right\rfloor,\\ l^{\prime}=1,\dots,\left\lfloor\frac{k}{2}\right\rfloor\end{array}\right\}.$ (101)
Proof ..
By the similar calculation as in the proof of Proposition 3.5, for $k\geq 2$, we have
$\displaystyle\mathrm{I}_{h}(\omega_{4}^{k-1}\omega_{0})$ $\displaystyle=\frac{1}{\sqrt{3}^{k-1}}\frac{1}{(k-1)!}\left\{\left(\frac{\pi}{3}\right)^{k-1}\log\frac{\sqrt{3}}{2}-\int_{0}^{\frac{\pi}{3}}t^{k-2}\log\sin t\leavevmode\nobreak\ dt\right\}$ (102)
$\displaystyle=\frac{1}{\sqrt{3}^{k-1}}\frac{1}{(k-1)!}\left\{\left(\frac{\pi}{3}\right)^{k-1}\log\frac{\sqrt{3}}{2}+(-1)^{k}(k-2)!(\mathrm{Li}_{k}(1)+\mathrm{Li}_{k}(-1))\right.$ (103)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\sum_{j=2}^{k}(-1)^{j-1}\frac{(k-1)!}{(k-j)!}\left(\frac{\pi\sqrt{-1}}{3}\right)^{k-j}\left(\mathrm{Li}_{j}(\xi_{6})-\frac{\pi\sqrt{-1}}{2}\frac{1}{(j-1)!}\left(\frac{\pi\sqrt{-1}}{3}\right)^{j-1}+\mathrm{Li}_{j}(-\xi_{6})\right)$ (104)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left.-\frac{\log 2}{k-1}\left(\frac{\pi\sqrt{-1}}{3}\right)^{k-1}+\frac{1}{k-1}\frac{\pi\sqrt{-1}}{2}\left(\frac{\pi\sqrt{-1}}{3}\right)^{k-1}-\frac{1}{k}\left(\frac{\pi\sqrt{-1}}{3}\right)^{k}\right\}.$ (105)
Here, for each $j\geq 2$,
$\displaystyle\mathrm{Li}_{j}(\xi_{6})+\mathrm{Li}_{j}(-\xi_{6})$ $\displaystyle=\sum_{n\geq 1}\frac{\xi_{6}^{n}+(-\xi_{6})^{n}}{n^{j}}$ (106)
$\displaystyle=2^{1-j}\sum_{n\geq 1}\frac{\xi_{3}^{n}}{n^{j}}$ (107)
$\displaystyle=2^{-j}(3^{1-j}-1)\zeta(j)+2^{-j}\sqrt{-3}L(j,\chi_{-3}).$ (108)
Then, the real part of the equation 105 is
$\displaystyle\frac{1}{\sqrt{3}^{k-1}}\frac{1}{(k-1)!}\left(\frac{\pi}{3}\right)^{k-1}\frac{1}{2}\log 3+\frac{1-(-1)^{k}}{2}(-1)^{\frac{1-k}{2}}\frac{1}{\sqrt{3}^{k-1}}2^{1-k}\zeta(k)$ (109)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{\sqrt{3}^{k-2}}\sum_{j=2\text{: even}}^{k}\frac{(-1)^{\frac{2-j}{2}}}{(k-j)!}\left(\frac{\pi}{3}\right)^{k-j}2^{-j}L(j,\chi_{-3})-\frac{1}{\sqrt{3}^{k-1}}\sum_{j=3\text{: odd}}^{k}\frac{(-1)^{\frac{1-j}{2}}}{(k-j)!}\left(\frac{\pi}{3}\right)^{k-j}2^{-j}(3^{1-j}-1)\zeta(j)$ (110)
and the equations 91 and 94 are shown by setting $k=2m$ and $k=2m+1$. The latter claims are also shown in the same way as the proof of Proposition 3.5. ∎
Example 3.1-(2), (3) and Proposition 3.5 suggest that $\mathrm{I}_{g}$ can be related to multiple $L$-values of level 4 and the parity of the powers of $\pi$ can be related to the parity of the number of $\omega_{2}$. Similarly, Proposition 3.7 suggests that $\mathrm{I}_{h}$ can be related to multiple $L$-values of level 6 and the parity of the powers of $\pi$ can be related to the parity of the number of $\omega_{4}$. These observations can be understood through Section 5.
4. Motivic Backgrounds
Here, we review some properties of the motivic iterated integrals studied by Deligne, Glanois, and Goncharov.
4.1. Motivic iterated integrals
Let $\overline{\mathbb{Q}}$ be an algebraic closure of $\mathbb{Q}$ and $F(\subseteq\overline{\mathbb{Q}})$ be a number field. Then, then there are a graded commutative Hopf algebra $\mathcal{A}_{F}$ and a graded commutative $\mathcal{A}_{F}$-comodule $\mathcal{H}_{F}$. We investigate the motivic structure of periods in these algebras. Here, we review some properties of them and define motivic iterated integrals.
Theorem 4.1 ([8, p.211, section 1.1]).
Let $\mathcal{A}_{F}$ be the fundamental Hopf algebra of the abelian category $\mathrm{MTM}_{F}$ of mixed Tate motives over $F$. Put $\mathcal{H}_{F}:=\mathcal{A}_{F}\otimes_{\mathbb{Q}}\mathbb{Q}[\tau]$ where $\tau$ is a formal element. Then, $\mathcal{A}_{F}$ and $\mathcal{H}_{F}$ satisfy following conditions.
  1. (1)
    There exists a $\mathbb{Q}$-algebra homomorphism $\mathrm{per}:\mathcal{H}_{F}\rightarrow\mathbb{C}$ such that $\mathrm{per}(\tau)=2\pi\sqrt{-1}$.
  2. (2)
    $\mathcal{A}_{F}$ is a graded commutative Hopf algebra. Let $\mathcal{A}_{F}^{(k)}$ be the $k$-th degree part of $\mathcal{A}_{F}$. In particular, there exists a $\mathbb{Q}$-linear homomorphism
    $\Delta:\mathcal{A}_{F}\rightarrow\mathcal{A}_{F}\otimes_{\mathbb{Q}}\mathcal{A}_{F},$
    which is called the coproduct on $\mathcal{A}_{F}$.
  3. (3)
    $\mathcal{H}_{F}$ is a graded commutative $\mathcal{A}$-comodule. Let $\mathcal{H}_{F}^{(k)}$ be the $k$-th degree part of $\mathcal{H}_{F}$. In particular, there exists a $\mathbb{Q}$-linear homomorphism
    $\Delta:\mathcal{H}_{F}\rightarrow\mathcal{A}_{F}\otimes_{\mathbb{Q}}\mathcal{H}_{F},$
    which is called the coaction on $\mathcal{H}_{F}$.
  4. (4)
    $\mathcal{A}_{F}^{(0)}=\mathcal{H}_{F}^{(0)}=\mathbb{Q}$.
  5. (5)
    There exists a $\mathbb{Q}$-algebra isomorphism $\mathcal{A}_{F}\cong\mathcal{H}_{F}/(\tau\cdot\mathcal{H}_{F})$. Let $\rho:\mathcal{H}_{F}\rightarrow\mathcal{A}_{F}$ be the natural projection.
Further, these objects satisfy some reasonable properties explained later.
Remark 4.2.
The period map $\mathrm{per}:\mathcal{H}_{F}\rightarrow\mathbb{C}$ is conjecturally injective (called period conjecture, see [5, p.342, Conjecture 4.118]). This conjecture suggests that information about the structure of the periods is not lose by the motivic lifting.
The motivic iterated integrals yield elements in $\mathcal{H}_{F}$.
Theorem 4.3 (Motivic Iterated Integrals, [8, p.212, Theorem 1.1]).
For each positive integer $k\in\mathbb{Z}_{\geq{0}}$, $k+2$ points $a_{0},\dots,a_{k+1}\in F$, two tangential base points $p=(a_{0},u),q=(a_{k+1},v)$, and a smooth path $\gamma$ from $p$ to $q$, there exists an element
$\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;a_{1},\dots,a_{k};q)\in\mathcal{H}_{F}^{(k)}$
called the motivic iterated integral, such that
$\mathrm{per}(\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;a_{1},\dots,a_{k};q))=\mathrm{I}_{\gamma}(p;a_{1},\dots,a_{k};q).$
Furthermore, the motivic iterated integrals satisfy following conditions.
  1. (1)
    The element $\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;a_{1},\dots,a_{k};q)$ is independent of the parametrization of $\gamma$.
  2. (2)
    The element $\rho(\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;a_{1},\dots,a_{k};q))$ is independent of the homotopy class of $\gamma$. We put
    $\mathrm{I}^{\mathfrak{a}}(p;a_{1},\dots,a_{k};q):=\rho(\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;a_{1},\dots,a_{k};q))\in\mathcal{A}_{N}^{(k)}.$
  3. (3)
    If $k=0$, $\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;q)=1$.
  4. (4)
    If $k>0$ and $p=q$, $\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;a_{1},\dots,a_{k};p)=0$.
  5. (5)
    The equations $\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;0;1)=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;1;1)=0$ hold.
  6. (6)
    (path reversal): For a smooth path $\gamma$ on $\mathbb{C}$,
    $\mathrm{I}^{\mathfrak{m}}_{\gamma^{-1}}(p;a_{1},\dots,a_{k};q)=(-1)^{k}\mathrm{I}^{\mathfrak{m}}_{\gamma}(q;a_{k},\dots,a_{1};p).$
  7. (7)
    (path connection): For smooth paths $\gamma_{1},\gamma_{2}$ on $\mathbb{C}$,
    $\mathrm{I}^{\mathfrak{m}}_{\gamma_{1}\gamma_{2}}(p;a_{1},\dots,a_{k};q)=\sum_{s=1}^{k}\mathrm{I}^{\mathfrak{m}}_{\gamma_{1}}(p;a_{1},\dots,a_{s};\gamma_{1}(1))\mathrm{I}^{\mathfrak{m}}_{\gamma_{2}}(\gamma_{2}(0);a_{s+1},\dots,a_{k};q).$
Definition 4.4.
For an element $w=e_{a_{1}}\dots e_{a_{k}}$ of non-commutative polynomial ring $W_{F}:=\mathbb{Q}\langle e_{a}\leavevmode\nobreak\ |\leavevmode\nobreak\ a\in F\rangle$, the motivic iterated integral $\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;w;q)\in\mathcal{H}_{F}$ is defined by
$\displaystyle\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;w;q):=\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;a_{1},\dots,a_{k};q),$ (111)
and extend to whole $W_{F}$ by $\mathbb{Q}$-linearity.
From here, we investigate the structures of $\mathcal{A}_{F}$ and $\mathcal{H}_{F}$ more precisely. For a $\mathbb{Q}$-subspace $R\subseteq F^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$, let $\mathcal{A}_{F,R}$ be a fundamental Hopf algebra of $R$. Then, it can be embedded into $\mathcal{A}_{F}$ (see [8, p.279, Appendix A.4]). We put $\mathcal{H}_{F,R}:=\mathcal{A}_{F,R}\otimes_{\mathbb{Q}}\mathbb{Q}[\tau]$.
Theorem 4.5 ([8, p.215, Theorem 1.3]).
Let $a_{0},\dots,a_{k+1}$ be elements of $F$. If $\widetilde{\mathrm{I}}(a_{i_{1}},a_{i_{2}},a_{i_{3}})\in R$ for any $1\displaystyle 0\leq i_{1}<i_{2}<i_{3}\leq k+$ 1, then it holds
$\displaystyle\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}_{a_{0},a_{k+1}}}(a_{0};a_{1},\dots,a_{k};a_{k+1})\in\mathcal{H}_{F,R}^{(k)}.$ (112)
By taking $p=0$, $q=1$ and $a_{1},\dots,a_{k}\in\widetilde{\mu}_{N}$, we have a motivic interpretation of multiple $L$-value.
Corollary 4.6.
For each $N\in\mathbb{Z}_{\geq{1}}$, $k_{1},\dots,k_{d}\in\mathbb{Z}_{\geq{1}}$ and $\alpha_{1},\dots,\alpha_{d}\in\mu_{N}$ with $(k_{d},\alpha_{d})\neq(1,1)$, there exits $L^{\mathfrak{m}}\left(\begin{matrix}k_{1},\dots,k_{d}\\ \alpha_{1},\dots,\alpha_{d}\end{matrix}\right)\in\mathcal{H}^{(k)}_{\mathbb{Q}(\xi_{N}),R_{N}}$ such that
$\displaystyle\mathrm{per}\left(L^{\mathfrak{m}}\left(\begin{matrix}k_{1},\dots,k_{d}\\ \alpha_{1},\dots,\alpha_{d}\end{matrix}\right)\right)=L\left(\begin{matrix}k_{1},\dots,k_{d}\\ \alpha_{1},\dots,\alpha_{d}\end{matrix}\right).$ (117)
Here, $R_{N}:=\langle 1-\xi_{N}^{r}\leavevmode\nobreak\ |\leavevmode\nobreak\ r=1,\dots,N-1\rangle_{\mathbb{Q}}\subseteq\mathbb{Q}(\xi_{N})^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$.
The Hopf structure of $\mathcal{H}_{F,R}$ is given by the following proposition.
Proposition 4.7 ([8, p.279, Appendix A.4 and A.5]).
There exists a non-canonical isomorphism
$\displaystyle\Phi:\mathcal{A}_{F,R}\overset{\sim}{\longrightarrow}T\left(R\oplus\bigoplus_{n=2}^{\infty}K_{2n-1}(F)\right),$ (118)
where $K_{2n-1}$ is in the degree $1\displaystyle 2n-$ 1 Quillen $K$-theory. In particular, there is a non-canonical isomorphism of Hopf algebras
$\displaystyle\Phi:\mathcal{A}_{F,R}\overset{\sim}{\longrightarrow}\mathbb{Q}\left\langle b_{j}^{(s)}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j\in\mathbb{Z}_{\geq{1}}\\ 1\leq s\leq s_{j}\end{array}\right\rangle=:\mathcal{V}_{F,R},$ (121)
where $\mathcal{V}_{F,R}$ is the non-commutative polynomial ring spanned by degree $j\geq 1$ elements $b_{j}^{s}$ and
$s_{j}:=\begin{cases}\dim_{\mathbb{Q}}R,&j=1,\\ r_{1}(F)+r_{2}(F),&j>1\text{ and }j\text{: odd},\\ r_{2}(F),&j\text{: even}.\end{cases}$
Also, there is an isomorphism
$\displaystyle\Phi:\mathcal{H}_{F,R}\overset{\sim}{\longrightarrow}\mathcal{V}_{F,R}\otimes_{\mathbb{Q}}\mathbb{Q}[\tau]=:\mathcal{U}_{F,R},$ (122)
Corollary 4.8.
Set $d_{k}:=\dim_{\mathbb{Q}}\mathcal{A}_{F,R}^{(k)}$. If $\dim_{\mathbb{Q}}R<\infty$, then
$\displaystyle\sum_{k=0}^{\infty}d_{k}t^{k}=\left(1-(\dim_{\mathbb{Q}}R)\cdot t-\frac{r_{2}t^{2}+(r_{1}+r_{2})t^{3}}{1-t^{2}}\right)^{-1},$ (123)
where $r_{1}:=r_{1}(F)$ and $r_{2}:=r_{2}(F)$. Also, set $d_{k}^{\prime}:=\dim_{\mathbb{Q}}\mathcal{H}_{F,R}^{(k)}$, then
$\displaystyle\sum_{k=0}^{\infty}d_{k}t^{k}=\left(1-(\dim_{\mathbb{Q}}R)\cdot t-\frac{r_{2}t^{2}+(r_{1}+r_{2})t^{3}}{1-t^{2}}\right)^{-1}\cdot(1-t)^{-1}.$ (124)
4.2. Specific cases
In this subsection, we review some properties of the structure of $\mathcal{H}_{F,R}$ for some specific cases used in Section 5. We consider the two cases: (I) $(F,R)=(\mathbb{Q}(\xi_{4}),\langle 2\rangle_{\mathbb{Q}})$, which relates to multiple $L$-values of level 4; (II) $(F,R)=(\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{\mathbb{Q}})$, which relates to multiple $L$-values of level 6.
Proposition 4.9 ([3, p.46, Théorème 8.9]).
The following conditions hold.
  1. (I)
    The $\mathbb{Q}$-linear space $\mathcal{H}_{\mathbb{Q}(\xi_{4}),\langle 2\rangle_{\mathbb{Q}}}^{(k)}$ is spanned by motivic iterated integrals formed
    $\displaystyle\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;a_{1},\dots,a_{k};1)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (a_{1},\dots,a_{k}\in\widetilde{\mu}_{4}).$ (125)
  2. (II)
    The $\mathbb{Q}$-linear space $\mathcal{H}_{\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{\mathbb{Q}}}^{(k)}$ is spanned by motivic iterated integrals formed
    $\displaystyle\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;a_{1},\dots,a_{k};1)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (a_{1},\dots,a_{k}\in\widetilde{\mu}_{6}).$ (126)
Proposition 4.10 ([4, p.49, Théorème 6.8]).
The following conditions hold.
  1. (I)
    For $k\in\mathbb{Z}_{\geq{1}}$, the elements formed $\mathrm{I}^{\mathfrak{a}}(0;\delta,\{0\}^{k-1};1)$ $(\delta\in\mu_{4})$ are subject only to the following relations in $\mathcal{A}_{\mathbb{Q}(\xi_{4}),\langle 2\rangle_{\mathbb{Q}}}$:
    $\displaystyle\begin{cases}\mathrm{I}^{\mathfrak{a}}(0;\delta,\{0\}^{k-1};1)=2^{k-1}\sum_{\epsilon^{2}=\delta}\mathrm{I}^{\mathfrak{a}}(0;\epsilon,\{0\}^{k-1};1),&\forall\delta\in\mu_{2},\\ \mathrm{I}^{\mathfrak{a}}(0;\delta,\{0\}^{k-1};1)=(-1)^{k-1}\mathrm{I}^{\mathfrak{a}}(0;\delta^{-1},\{0\}^{k-1};1),&\forall\delta\in\mu_{4}.\end{cases}$ (127)
  2. (II)
    For $k\in\mathbb{Z}_{\geq{1}}$, the elements formed $\mathrm{I}^{\mathfrak{a}}(0;\delta,\{0\}^{k-1};1)$ $(\delta\in\mu_{6})$ are subject only to the following relations in $\mathcal{A}_{\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{\mathbb{Q}}}$:
    $\displaystyle\begin{cases}\mathrm{I}^{\mathfrak{a}}(0;\delta,\{0\}^{k-1};1)=2^{k-1}\sum_{\varepsilon^{2}=\delta}\mathrm{I}^{\mathfrak{a}}(0;\varepsilon,\{0\}^{k-1};1),&\forall\delta\in\mu_{3},\\ \mathrm{I}^{\mathfrak{a}}(0;\delta,\{0\}^{k-1};1)=3^{k-1}\sum_{\varepsilon^{3}=\delta}\mathrm{I}^{\mathfrak{a}}(0;\varepsilon,\{0\}^{k-1};1),&\forall\delta\in\mu_{2},\\ \mathrm{I}^{\mathfrak{a}}(0;\delta,\{0\}^{k-1};1)=(-1)^{k-1}\mathrm{I}^{\mathfrak{a}}(0;\delta^{-1},\{0\}^{k-1};1)&\forall\delta\in\mu_{6}.\end{cases}$ (128)
Proposition 4.11 ([6, p.354, Corollary 3.7]).
We put $\Delta^{\prime}:=\Delta-1\otimes\mathrm{id}:\mathcal{H}_{F,R}\rightarrow\mathcal{A}_{F,R}\otimes_{\mathbb{Q}}\mathcal{H}_{F,R}$. Then, the following equations hold:
$\displaystyle\mathrm{Ker}\Delta^{\prime}\cap\mathcal{H}_{\mathbb{Q}(\xi_{4}),\langle 2\rangle_{\mathbb{Q}}}^{(k)}=\mathbb{Q}\tau^{k}\oplus\mathbb{Q}\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{4},\{0\}^{k-1};1);$ (129)
$\displaystyle\mathrm{Ker}\Delta^{\prime}\cap\mathcal{H}_{\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{\mathbb{Q}}}^{(k)}=\mathbb{Q}\tau^{k}\oplus\mathbb{Q}\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{6},\{0\}^{k-1};1).$ (130)
5. Main Results
5.1. Base extension
First, we consider the pullback of $B_{f}$ under $\phi_{f}:\mathbb{P}^{1}(\mathbb{C})\rightarrow X_{f}$ to use Section 4 for giving a motivic interpretation of $\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})$.
Pullback of $B_{f}$.
  1. (I)
    $g(X,Y,Z)=X^{2}+Y^{2}-Z^{2}$: for each element of $B_{g}=\left\{\omega_{0}:=\frac{\leavevmode\nobreak\ dx}{x},\omega_{1}:=\frac{\leavevmode\nobreak\ dx}{1-x},\omega_{2}:=\frac{\leavevmode\nobreak\ dx}{y},\omega_{3}:=\frac{\leavevmode\nobreak\ dx}{xy}\right\}$, the pullback along $\phi_{g}$ is
    $\displaystyle\begin{split}\phi_{g}^{*}\omega_{0}&=\frac{(1-\lambda^{2})\leavevmode\nobreak\ d\lambda}{\lambda(1+\lambda^{2})}=\frac{\leavevmode\nobreak\ d\lambda}{\lambda}-\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{4}}-\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{4}^{-1}},\\ \phi_{g}^{*}\omega_{1}&=\frac{2(1+\lambda)\leavevmode\nobreak\ d\lambda}{(1-\lambda)(1+\lambda^{2})}=-2\frac{\leavevmode\nobreak\ d\lambda}{1-\lambda}+\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{4}}+\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{4}^{-1}},\\ \phi_{g}^{*}\omega_{2}&=\frac{2\leavevmode\nobreak\ d\lambda}{1+\lambda^{2}}=\xi_{4}^{-1}\left(\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{4}}-\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{4}^{-1}}\right),\\ \phi_{g}^{*}\omega_{3}&=\frac{\leavevmode\nobreak\ d\lambda}{\lambda}.\end{split}$
    The integral path is $\alpha:[0,1]\rightarrow\mathbb{P}^{1}(\mathbb{C})$, $\alpha(t)=1-\sqrt{\frac{1-t}{1+t}}$ and it satisfies $\alpha([0,1])=[0,1]$. Then, each $\mathrm{I}_{g}(\eta_{1}\cdots\eta_{k})$ can be written by a $\mathbb{Q}(\xi_{4})$-linear combination of iterated integrals on $\mathbb{P}^{1}(\mathbb{C})\setminus\{0,1,\pm\sqrt{1},\infty\}$ formed
    $\displaystyle\mathrm{I}_{\mathrm{dch}}(0;a_{1},\dots,a_{k};1)$ (131)
    with $a_{j}\in F_{g}:=\mathbb{Q}(\xi_{4})$ for $j=1,\dots,k$. Also, for each of the terms formed as the equation 131, it holds $\widetilde{\mathrm{I}}(a_{i_{1}},a_{i_{2}},a_{i_{3}})\in\langle 2\rangle_{\mathbb{Q}}=:R_{g}\subseteq F_{g}^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ if $1\displaystyle 0\leq i_{1}<i_{2}<i_{3}\leq k+$ 1.
  2. (II)
    $h(X,Y,Z)=X^{2}+XY+Y^{2}-Z^{2}$: for each element of $B_{h}=\left\{\omega_{0}:=\frac{\leavevmode\nobreak\ dx}{x},\omega_{1}:=\frac{\leavevmode\nobreak\ dx}{1-x},\omega_{4}:=\frac{\leavevmode\nobreak\ dx}{x+2y},\right.$
    $\left.\omega_{5}:=\frac{\leavevmode\nobreak\ dx}{x(x+2y)},\omega_{6}:=\frac{\leavevmode\nobreak\ dx}{(1-x)(x+2y)}\right\}$, the pullback along $\phi_{h}$ is
    $\displaystyle\phi_{h}^{*}\omega_{0}$ $\displaystyle=\frac{-(\lambda^{2}-2\lambda-2)\leavevmode\nobreak\ d\lambda}{\lambda(2+\lambda)(1+\lambda+\lambda^{2})}=\frac{\leavevmode\nobreak\ d\lambda}{\lambda}+\frac{\leavevmode\nobreak\ d\lambda}{\lambda+2}-\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{3}}-\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{3}^{-1}},$ (132)
    $\displaystyle\phi_{h}^{*}\omega_{1}$ $\displaystyle=\frac{(\lambda^{2}-2\lambda-2)\leavevmode\nobreak\ d\lambda}{\lambda^{3}-1}=-\frac{\leavevmode\nobreak\ d\lambda}{\lambda-1}+\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{3}}+\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{3}^{-1}},$ (133)
    $\displaystyle\phi_{h}^{*}\omega_{4}$ $\displaystyle=\frac{\leavevmode\nobreak\ d\lambda}{1+\lambda+\lambda^{2}}=\frac{1}{\sqrt{-3}}\left(\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{3}}-\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\xi_{3}^{-1}}\right),$ (134)
    $\displaystyle\phi_{h}^{*}\omega_{5}$ $\displaystyle=\frac{\leavevmode\nobreak\ d\lambda}{\lambda(\lambda+2)}=\frac{1}{2}\left(\frac{\leavevmode\nobreak\ d\lambda}{\lambda}-\frac{\leavevmode\nobreak\ d\lambda}{\lambda+2}\right),$ (135)
    $\displaystyle\phi_{h}^{*}\omega_{6}$ $\displaystyle=-\frac{\leavevmode\nobreak\ d\lambda}{\lambda-1}.$ (136)
    The integral path is $\alpha:[0,1]\rightarrow\mathbb{P}^{1}(\mathbb{C})$, $\alpha(t)=\frac{t-2+\sqrt{4-3t^{2}}}{2(1-t)}$ and it satisfies $[0,1]$. Then, each $\mathrm{I}_{h}(\eta_{1}\cdots\eta_{k})$ can be written by a $\mathbb{Q}(\xi_{6})$-linear combination of iterated integrals on $\mathbb{P}^{1}(\mathbb{C})\setminus\{0,1,-2,\xi_{3}^{\pm 1},\infty\}$ formed
    $\displaystyle\mathrm{I}_{\mathrm{dch}}(0;a_{1},\dots,a_{k};1)$ (137)
    with $a_{j}\in F_{h}:=\mathbb{Q}(\xi_{6})$ for $j=1,\dots,k$. Also, for each of the terms formed as the equation 137, it holds $\widetilde{\mathrm{I}}(a_{i_{1}},a_{i_{2}},a_{i_{3}})\in\langle 2,3\rangle_{\mathbb{Q}}=:R_{h}\subseteq F_{h}^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ if $1\displaystyle 0\leq i_{1}<i_{2}<i_{3}\leq k+$ 1.
By the above discussion, for each $f\in\{g,h\}$ and $(\eta_{1},\dots,\eta_{k})\in B_{f}^{(k)}$, $\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})$ can be written as an $F_{f}$-linear combination of iterated integrals formed
$\displaystyle\mathrm{I}_{\mathrm{dch}}(0;a_{1},\dots,a_{k};1)$ (138)
with $a_{j}\in F_{f}$ for $j=1,\dots,k$. We review that the $\mathbb{Q}$-linear combination of the equation 138 has a motivic interpretation given by motivic iterated integrals in $\mathcal{H}_{F_{f},R_{f}}$. Then, we consider the base extension of $\mathcal{H}_{F_{f},R_{f}}$ to give motivic interpretations of iterated integrals on $X_{f}$.
Definition 5.1.
For a number field $F\supseteq\mathbb{Q}$ and $R\subseteq F^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$, we extend the following notations to $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$. For $u\otimes c,u_{1}\otimes c_{1},u_{2}\otimes c_{2}\in\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$,
  • coaction: $\Delta:\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F\rightarrow\mathcal{A}_{F,R}\otimes_{\mathbb{Q}}\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F,\leavevmode\nobreak\ \leavevmode\nobreak\ \Delta(u\otimes c):=\Delta(u)\otimes c$;
  • product: $(\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F)\otimes_{\mathbb{Q}}(\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F)\rightarrow\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F,\leavevmode\nobreak\ \leavevmode\nobreak\ (u_{1}\otimes c_{1})\cdot(u_{2}\otimes c_{2}):=u_{1}u_{2}\otimes c_{1}c_{2}$;
  • period map: $\mathrm{per}:\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F\rightarrow\mathbb{C},\leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{per}(u\otimes c):=\mathrm{per}(u)\cdot c$;
  • isomorphism: $\Phi:\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F\rightarrow\mathcal{U}_{F,R}\otimes_{\mathbb{Q}}F,\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi(u\otimes c):=\Phi(u)\otimes c$.
Remark 5.2.
As same as [14, p.19, Proposition 4.4.2], if the period conjecture is true, (i.e. the period map $\mathrm{per}:\mathcal{H}_{F,R}\rightarrow\mathbb{C}$ is injective), then its extension $\mathrm{per}:\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F\rightarrow\mathbb{C}$ is also injective. This suggests that information about the structure of the periods is not also lost by the base extension.
Definition 5.3.
Let $w=c\cdot e_{a_{1}}\dots e_{a_{k}}$ be an element of non-commutative polynomial ring $F\langle F\rangle$ with $c\in F$, $a_{1},\dots,a_{k}\in F$ satisfying $\widetilde{\mathrm{I}}(a_{p},a_{q},a_{r})\in R$. Then, for a smooth path $\gamma:[0,1]\rightarrow\mathbb{C}$ connecting $p\in\mathbb{C}$ and $q\in\mathbb{C}$, the motivic iterated integral $\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;w;q)\in\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ is defined by
$\displaystyle\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;w;q):=\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;a_{1},\dots,a_{k};q)\otimes c.$ (139)
We also extend $\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;w;q)$ $F$-linearity to whole $w\in F\langle F\rangle$.
Motivic interpretations of iterated integrals on $X_{f}$ are given as elements of $\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}$.
Definition 5.4.
For $f\in\{g,h\}$, we define the motivic iterated integrals on $X_{f}$ as follows.
  1. (I)
    For $(\eta_{1},\dots,\eta_{k})\in B_{g}^{(k)}$, we define
    $\displaystyle\mathrm{I}^{\mathfrak{m}}_{g}(\eta_{1}\cdots\eta_{k}):=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\theta_{1}\cdots\theta_{k};1)\in\mathcal{H}_{F_{g},R_{g}}\otimes_{\mathbb{Q}}F_{g}$ (140)
    where $\theta_{1},\cdots,\theta_{k}\in F_{g}\langle e_{a}\leavevmode\nobreak\ |\leavevmode\nobreak\ a\in F_{g}\rangle$ are defined by
    $\displaystyle\begin{split}\theta_{j}:=\begin{cases}e_{0}-e_{\xi_{4}}-e_{\xi_{4}^{-1}},&\eta_{j}=\omega_{0},\\ -2e_{1}+e_{\xi_{4}}+e_{\xi_{4}^{-1}},&\eta_{j}=\omega_{1},\\ \xi_{4}^{-1}(e_{\xi_{4}}-e_{\xi_{4}^{-1}}),&\eta_{j}=\omega_{2},\\ e_{0},&\eta_{j}=\omega_{3}.\end{cases}\end{split}$
  2. (II)
    For $(\eta_{1},\dots,\eta_{k})\in B_{h}^{(k)}$, we define
    $\displaystyle\mathrm{I}^{\mathfrak{m}}_{h}(\eta_{1}\cdots\eta_{k}):=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\theta_{1}\cdots\theta_{k};1)\in\mathcal{H}_{F_{h},R_{h}}\otimes_{\mathbb{Q}}F_{h}$ (141)
    where $\theta_{1},\cdots,\theta_{k}\in F_{h}\langle e_{a}\leavevmode\nobreak\ |\leavevmode\nobreak\ a\in F_{h}\rangle$ are defined by
    $\displaystyle\theta_{j}:=\begin{cases}e_{0}+e_{-2}-e_{\xi_{3}}-e_{\xi_{3}^{-1}},&\eta_{j}=\omega_{0},\\ -e_{1}+e_{\xi_{3}}+e_{\xi_{3}^{-1}},&\eta_{j}=\omega_{1},\\ \frac{1}{\sqrt{-3}}(e_{\xi_{3}}-e_{\xi_{3}^{-1}}),&\eta_{j}=\omega_{4},\\ \frac{1}{2}(e_{0}-e_{-2}),&\eta_{j}=\omega_{5},\\ -e_{1},&\eta_{j}=\omega_{6}.\end{cases}$ (142)
Then, by Definition 5.3 and the pullback of $B_{f}$, we have
$\displaystyle\mathrm{per}(\mathrm{I}^{\mathfrak{m}}_{f}(\eta_{1}\cdots\eta_{k}))=\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})$ (143)
for each $f\in\{g,h\}$ and $(\eta_{1},\dots,\eta_{k})\in B_{f}^{(k)}$. This allows us to give a motivic interpretation of the iterated integrals on $X_{f}$.
5.2. Galois action
Now, we consider the action of the Galois group $\mathrm{Gal}(F/\mathbb{Q})$ on $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ to investigate the structure of $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ in more detail.
The Galois group $\mathrm{Gal}(F/\mathbb{Q})$ acts on the category of mixed Tate motives $\mathrm{MTM}_{R}$ on $R\subseteq F^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ induced by the automorphism as a ring of $R$ induced by the Galois action. In addition, the Galois action is induced on $\mathcal{H}_{F,R}$ since $\mathcal{H}_{F,R}$ is the coordinate ring of Tannakian fundamental group of $\mathrm{MTM}_{R}$. Also, we can extend the $\mathrm{Gal}(F/\mathbb{Q})$-action to $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ diagonally. By Proposition 4.9, each element of $\mathcal{H}_{F,R}$ can be written by motivic iterated integrals if $(F,R)=(F_{g},R_{g})$ or $(F,R)=(F_{h},R_{h})$, then we define the Galois action explicitly for motivic iterated integrals.
Proposition 5.5.
If $(F,R)=(F_{f},R_{f})$ for $f\in\{g,h\}$, the Galois group $\mathrm{Gal}(F/\mathbb{Q})$ acts on $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$. In addition, $\sigma\in\mathrm{Gal}(F/\mathbb{Q})$ acts on the element written by motivic iterated integrals as
$\displaystyle\sigma(\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;a_{1},\dots,a_{k};1)\otimes c)=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;{}^{\sigma}{a_{1}},\dots,{}^{\sigma}{a_{k}};1)\otimes{}^{\sigma}{c}.$ (144)
Let us look back at Definition 5.4. Recall that we put $G_{f}:=\mathrm{Gal}(F_{f}/\mathbb{Q})$. Then, we see that each element $\mathrm{I}^{\mathfrak{m}}_{f}(\eta_{1}\cdots\eta_{k})\in\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}$ is $G_{f}$-invariant since so is each $\theta_{p}$ $(p=1,\dots,k)$. Then, it is sufficient to investigate the $G_{f}$-invariant part $\left(\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}$ of $\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}$ to discuss the motivic structure of iterated integrals on $X_{f}$.
Theorem 5.6 (Theorem 1.2).
For each $f\in\{g,h\}$ and $(\eta_{1},\dots,\eta_{k})\in B_{f}^{(k)}$, there exists an element
$\mathrm{I}^{\mathfrak{m}}_{f}(\eta_{1}\cdots\eta_{k})\in\left(\mathcal{H}_{F_{f},R_{f}}^{(k)}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}$
such that $\mathrm{per}(\mathrm{I}^{\mathfrak{m}}_{f}(\eta_{1}\cdots\eta_{k}))=\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})$.
5.3. The structure of $\left(\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}$
Now, we investigate the structure of $\left(\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}$ and reduce to the original iterated integrals $\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})$.
First, in preparation, we write down Proposition 4.7 explicitly for each case of $\mathcal{H}_{F_{f},R_{f}}$.
  1. (I)
    if $f=g$, we have
    $\displaystyle\mathcal{U}_{F_{g},R_{g}}=\mathbb{Q}\langle b_{j}\leavevmode\nobreak\ |\leavevmode\nobreak\ j\in\mathbb{Z}_{\geq{1}}\rangle\otimes_{\mathbb{Q}}\mathbb{Q}[\tau].$ (145)
    In particular, the degree $k$ part $\mathcal{U}_{F_{g},R_{g}}^{(k)}$ is
    $\displaystyle\bigoplus_{r\geq 0}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}\cdots b_{j_{r}}\cdot\tau^{l}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}.$ (148)
  2. (II)
    if $f=h$, we have
    $\displaystyle\mathcal{U}_{F_{h},R_{h}}=\mathbb{Q}\left\langle b_{j}^{(s)}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ j\in\mathbb{Z}_{\geq{1}},s\in\begin{cases}\{1,2\},&j=1\\ \{1\},&j>1\end{cases}\right\rangle\otimes_{\mathbb{Q}}\mathbb{Q}[\tau].$ (149)
    In particular, the degree $k$ part $\mathcal{U}_{F_{h},R_{h}}^{(k)}$ is
    $\displaystyle\bigoplus_{r\geq 0}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}^{(s_{1})}\cdots b_{j_{r}}^{(s_{r})}\cdot\tau^{l}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ s_{i}\in\begin{cases}\{1,2\},&j_{i}=1\\ \{1\},&j_{i}>1\end{cases},\leavevmode\nobreak\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}.$ (152)
Here, we fix one isomorphism given in Proposition 4.7. This allows us to write down the Galois action on $\mathcal{U}_{F_{f},R_{f}}$ induced by $\widetilde{\Phi}:\mathcal{H}_{F_{f},R_{f}}\rightarrow\mathcal{U}_{F_{f},R_{f}}$ concisely.
Proposition 5.7.
For each $f\in\{g,h\}$, we can normalize a non-canonical isomorphism $\Phi:\mathcal{H}_{F_{f},R_{f}}\rightarrow\mathcal{U}_{F_{f},R_{f}}$ in Proposition 4.7 as follows.
  1. (I)
    If $f=g$, the normalized isomorphism $\widetilde{\Phi}:\mathcal{H}_{F_{g},R_{g}}\rightarrow\mathcal{U}_{F_{g},R_{g}}$ is defined by
    $\displaystyle\widetilde{\Phi}(u_{k}):=b_{k},$ (153)
    where
    $\displaystyle u_{k}:=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{4},\{0\}^{k-1};1)-(-1)^{k}\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{4}^{-1},\{0\}^{k-1};1)\in\mathcal{H}_{F_{g},R_{g}}.$ (154)
  2. (II)
    If $f=h$, the normalized isomorphism $\widetilde{\Phi}:\mathcal{H}_{F_{h},R_{h}}\rightarrow\mathcal{U}_{F_{h},R_{h}}$ is defined by
    $\displaystyle\widetilde{\Phi}(v_{k}):=b_{k}^{(1)},$ (155)
    $\displaystyle\widetilde{\Phi}(v):=b_{1}^{(2)},$ (156)
    where
    $\displaystyle v_{k}:=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{3},\{0\}^{k-1};1)-(-1)^{k}\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{3}^{-1},\{0\}^{k-1};1)\in\mathcal{H}_{F_{h},R_{h}},$ (157)
    $\displaystyle v:=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;-1;1)\in\mathcal{H}_{F_{h},R_{h}}.$ (158)
Proof ..
(I) is shown in [14, p.21, Proposition 5.1.1]. Then, we only show (II). It is sufficient to prove that each of $\{v_{1},v,\tau\}$ and $\{v_{k},\tau^{k}\}\leavevmode\nobreak\ (k>1)$ is $\mathbb{Q}$-linear independent. First, by Proposition 4.10, we have
$\displaystyle\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{3},\{0\}^{k-1};1)+(-1)^{k}\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{3}^{-1},\{0\}^{k-1};1)\in\tau^{k}\mathbb{Q}$ (159)
for any $k\geq 1$. On the other hands, by Proposition 4.11, each of $\{v,\tau\}$ and $\{\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{3}^{-1},\{0\}^{k-1};1),\tau^{k}\}$ is $\mathbb{Q}$-linear independent. Thus, $\{v_{k},\tau^{k}\}$ is also $\mathbb{Q}$-linear independent. Furthermore, looking back at Proposition 4.10, there are no $\mathbb{Q}$-linear relations between $v$ and $v_{1}$. Then, $\{v,v_{1},\tau\}$ is $\mathbb{Q}$-linear independent.
Now, we calculate the action of the Galois group $G_{f}$ on $\mathcal{U}_{F_{f},R_{f}}$ induced by the isomorphisms $\widetilde{\Phi}:\mathcal{H}_{F_{f},R_{f}}\rightarrow\mathcal{U}_{F_{f},R_{f}}$. First, we calculate the action on each of the generators $b_{j}^{(s)}$ of $\mathcal{U}_{F_{f},R_{f}}$. Recall that $G_{f}:=\mathrm{Gal}(F_{f}/\mathbb{Q})$.
  1. (I)
    If $f=g$, $G_{g}=\{1,\sigma\}$ where $\sigma$ is defined by $\sigma(\xi_{4})=\xi_{4}^{-1}$. By definition of $u_{k}$, we have
    $\displaystyle{}^{\sigma}u_{k}=(-1)^{k+1}u_{k}.$ (160)
    Then, we have ${}^{\sigma}b_{k}=(-1)^{k+1}b_{k}$. Also, we have ${}^{\sigma}\tau=-\tau$.
  2. (II)
    If $f=h$, $G_{h}=\{1,\sigma\}$ where $\sigma$ is defined by $\sigma(\sqrt{-3})=-\sqrt{-3}$. By definition of $v$ and $v_{k}$, we have
    $\displaystyle{}^{\sigma}v=v,$ (161)
    $\displaystyle{}^{\sigma}v_{k}=(-1)^{k+1}v_{k}.$ (162)
    Then, we have ${}^{\sigma}b_{k}^{(1)}=b_{k}^{(1)}\leavevmode\nobreak\ (k\geq 1)$ and ${}^{\sigma}b_{1}^{(2)}=b_{1}^{(2)}$. Also, we have ${}^{\sigma}\tau=-\tau$.
The Galois action on $\mathcal{U}_{F,R}$ is commutative with the non-commutative product of $\mathcal{U}_{F,R}$. That fact is shown in the case of $(F,R)=(\mathbb{Q}(\xi_{4}),\langle 2\rangle_{\mathbb{Q}})$ in [14, p.23, Lemma 5.2.1] and we can show similarly if $(F,R)=(\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{\mathbb{Q}})$.
Lemma 5.8.
For the generators $b_{j_{p}}^{(s_{p})}\leavevmode\nobreak\ (p=1,\dots,r)$ of $\mathcal{U}_{F_{f},R_{f}}$, a non-negative integer $l\geq 0$, and $\sigma\in\mathrm{Gal}(F_{f}/\mathbb{Q})$ the following equation holds:
($\displaystyle{}^{\sigma}(b_{j_{1}}^{(s_{1})}\dots b_{j_{r}}^{(s_{r})}\cdot\tau^{l})={}^{\sigma}b_{j_{1}}^{(s_{1})}\dots{}^{\sigma}b_{j_{r}}^{(s_{r})}\cdot{}^{\sigma}(\tau^{l}).$ (163)
Then, we can identify the $G_{f}$-invariant part $\left(\mathcal{U}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}$ of $\mathcal{U}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}$.
Proposition 5.9.
The following conditions hold.
  1. (I)
    If $f=g$,
    $\displaystyle\left(\mathcal{U}_{F_{g},R_{g}}^{(k)}\otimes_{\mathbb{Q}}F_{g}\right)^{G_{g}}$ $\displaystyle=\bigoplus_{\begin{subarray}{c}r\geq 0\\ k-r\text{: even}\end{subarray}}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}\cdots b_{j_{r}}\cdot\tau^{l}\otimes 1\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}$ (168)
    $\displaystyle\oplus\bigoplus_{\begin{subarray}{c}r\geq 0\\ k-r\text{: odd}\end{subarray}}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}\cdots b_{j_{r}}\cdot\tau^{l}\otimes\sqrt{-1}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}.$ (173)
    In particular, it holds
    $\dim_{\mathbb{Q}}\left(\mathcal{U}_{F_{g},R_{g}}^{(k)}\otimes_{\mathbb{Q}}F_{g}\right)^{G_{g}}=d_{k,g},$
    where $\sum_{k=0}^{\infty}D_{g}^{(k)}t^{k}=\frac{1}{1-2t}$.
  2. (II)
    If $f=h$,
    $\displaystyle\left(\mathcal{U}_{F_{h},R_{h}}^{(k)}\otimes_{\mathbb{Q}}F_{h}\right)^{G_{h}}$ $\displaystyle=\bigoplus_{\begin{subarray}{c}r\geq 0\\ k-r\text{: even}\end{subarray}}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}^{(s_{1})}\cdots b_{j_{r}}^{(s_{r})}\cdot\tau^{l}\otimes 1\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ s_{i}\in\begin{cases}\{1,2\},&j_{i}=1\\ \{1\},&j_{i}>1\end{cases},\leavevmode\nobreak\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}$ (178)
    $\displaystyle\oplus\bigoplus_{\begin{subarray}{c}r\geq 0\\ k-r\text{: odd}\end{subarray}}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}^{(s_{1})}\cdots b_{j_{r}}^{(s_{r})}\cdot\tau^{l}\otimes\sqrt{-3}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ s_{i}\in\begin{cases}\{1,2\},&j_{i}=1\\ \{1\},&j_{i}>1\end{cases},\leavevmode\nobreak\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}.$ (183)
    In particular, it holds
    $\dim_{\mathbb{Q}}\left(\mathcal{U}_{F_{h},R_{h}}^{(k)}\otimes_{\mathbb{Q}}F_{d}\right)^{G_{h}}=D_{h}^{(k)},$
    where $\sum_{k=0}^{\infty}D_{h}^{(k)}t^{k}=\frac{1}{1-3t+t^{2}}$.
Here, we reduce them to $\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})\in\mathbb{C}$ by the period map $\mathrm{per}:\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}\rightarrow\mathbb{C}$. Since $Z_{f}\subseteq\mathrm{per}\left(\left(\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}\right)$, we have the following corollary.
Corollary 5.10 (Corollary 1.3).
For $f\in\{g,h\}$, it holds
$\displaystyle\dim_{\mathbb{Q}}Z_{f}^{(k)}\leq D_{f}^{(k)}.$ (184)
The following table summarizes the value of $D_{f}^{(k)}$ and conjectural value of $\dim_{\mathbb{Q}}Z_{f}^{(k)}$ for each $f\in\{g,h\}$ and small weight $k\in\mathbb{Z}_{\geq{1}}$ obtained by numerical experiments. The value of each iterated integral was calculated by SageMath, and linear independence was determined by the PARI/GP library of SageMath. As an observation of this table, we can check Corollary 1.3 numerically for $k=0,1,\dots,5$. The table also shows that there is rich relational $\mathbb{Q}$-linear relations along iterated integrals on $Y_{f}$. The difference between $D_{f}^{(k)}$ and $\dim_{\mathbb{Q}}Z_{f}^{(k)}$ is due to the fact that $\mathrm{per}\left(\left(\mathcal{H}_{F_{f},R_{f}}^{(k)}\otimes_{\mathbb{Q}}F_{f}\right)^{\mathrm{Gal}(F_{f}/\mathbb{Q})}\right)\not\subseteq Z_{f}^{(k)}$. Also, if $f=g$, $\dim_{\mathbb{Q}}Z_{g}^{(k)}$ is expected to be $1\displaystyle 2^{k}-$ 1 and the element $u\in\left(\mathcal{H}_{F_{f},R_{f}}^{(k)}\otimes_{\mathbb{Q}}F_{f}\right)^{\mathrm{Gal}(F_{f}/\mathbb{Q})}$ such that $\mathrm{per}(u)\notin Z_{g}^{(k)}$ would be $u=(\log^{\mathfrak{m}}2)^{k}:=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;-1;1)^{k}$ according to the numerical experiments.
Таблица 1. dimension evaluation
$k$ 00 0 11 11 22 22 33 33 44 44 55 55
#$\#B_{g}^{(k)}$ 11 11 66 66 2424 2424 9696 9696 384384 384384
$D_{g}^{(k)}$ 11 11 22 22 44 44 88 88 1616 1616 3232 3232
$\dim_{\mathbb{Q}}Z_{g}^{(k)}$? 11 11 11 11 33 33 77 77 1515 1515 3131 3131
#$\#B_{h}^{(k)}$ 11 11 99 99 4545 4545 225225 225225 11251125 11251125
$D_{h}^{(k)}$ 11 11 33 33 88 88 2121 2121 $5$ $5$ 144144 144144
$\dim_{\mathbb{Q}}Z_{h}^{(k)}$? 11 11 11 11 55 55 1515 1515 4646 4646 105105 105105
We also have the following corollary by considering where each $\mathrm{I}^{\mathfrak{m}}_{f}(\eta_{1}\cdots\eta_{k})$ belongs in the direct sum decomposition in Proposition 5.9.
Corollary 5.11.
The following conditions hold.
  1. (I)
    If $f=g$, we put
    $\displaystyle Z_{g,0}^{(k)}:=\left\langle\mathrm{I}_{g}(\eta_{1}\cdots\eta_{k})\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \#\left\{p\leavevmode\nobreak\ |\leavevmode\nobreak\ \eta_{p}=\omega_{2}\right\}\text{ is even}\right\rangle_{\mathbb{Q}},$ (185)
    $\displaystyle Z_{g,1}^{(k)}:=\left\langle\mathrm{I}_{g}(\eta_{1}\cdots\eta_{k})\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \#\left\{p\leavevmode\nobreak\ |\leavevmode\nobreak\ \eta_{p}=\omega_{2}\right\}\text{ is odd}\right\rangle_{\mathbb{Q}}.$ (186)
    Then,
    $\displaystyle Z_{g,0}^{(k)}\subseteq\mathrm{per}\left(\bigoplus_{\begin{subarray}{c}r\geq 0\\ k-r\text{: even}\end{subarray}}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}\cdots b_{j_{r}}\cdot\tau^{l}\otimes 1\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}\right),$ (191)
    $\displaystyle Z_{g,1}^{(k)}\subseteq\mathrm{per}\left(\bigoplus_{\begin{subarray}{c}r\geq 0\\ k-r\text{: odd}\end{subarray}}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}\cdots b_{j_{r}}\cdot\tau^{l}\otimes\sqrt{-1}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}\right).$ (196)
    In particular, for each $\delta\in\{0,1\}$, it holds
    $\displaystyle\dim_{\mathbb{Q}}Z_{g,\delta}^{(k)}\leq D_{g,\delta}^{(k)},$ (197)
    where
    $\displaystyle\sum_{k=0}^{\infty}D_{g,0}^{(k)}t^{k}=\frac{1-t}{1-2t},$ (198)
    $\displaystyle\sum_{k=0}^{\infty}D_{g,1}^{(k)}t^{k}=\frac{t}{1-2t}.$ (199)
  2. (II)
    If $f=h$, we put
    $\displaystyle Z_{h,0}^{(k)}:=\left\langle\mathrm{I}_{h}(\eta_{1}\cdots\eta_{k})\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \#\left\{i\leavevmode\nobreak\ |\leavevmode\nobreak\ \eta_{i}=\omega_{4}\right\}\text{ is even}\right\rangle_{\mathbb{Q}},$ (200)
    $\displaystyle Z_{h,1}^{(k)}:=\left\langle\mathrm{I}_{h}(\eta_{1}\cdots\eta_{k})\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \#\left\{i\leavevmode\nobreak\ |\leavevmode\nobreak\ \eta_{i}=\omega_{4}\right\}\text{ is odd}\right\rangle_{\mathbb{Q}}.$ (201)
    Then,
    $\displaystyle Z_{h,0}^{(k)}\subseteq\mathrm{per}\left(\bigoplus_{\begin{subarray}{c}r\geq 0\\ k-r\text{: even}\end{subarray}}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}^{(s_{1})}\cdots b_{j_{r}}^{(s_{r})}\cdot\tau^{l}\otimes 1\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ s_{i}\in\begin{cases}\{1,2\},&j_{i}=1\\ \{1\},&j_{i}>1\end{cases},\leavevmode\nobreak\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}\right),$ (206)
    $\displaystyle Z_{h,1}^{(k)}\subseteq\mathrm{per}\left(\bigoplus_{\begin{subarray}{c}r\geq 0\\ k-r\text{: odd}\end{subarray}}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}^{(s_{1})}\cdots b_{j_{r}}^{(s_{r})}\cdot\tau^{l}\otimes\sqrt{-3}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ s_{i}\in\begin{cases}\{1,2\},&j_{i}=1\\ \{1\},&j_{i}>1\end{cases},\leavevmode\nobreak\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}\right).$ (211)
    In particular, for each $\delta\in\{0,1\}$, it holds
    $\displaystyle\dim_{\mathbb{Q}}Z_{h,\delta}^{(k)}\leq D_{h,\delta}^{(k)},$ (212)
    where
    $\displaystyle\sum_{k=0}^{\infty}D_{h,0}^{(k)}t^{k}=\frac{1-t}{(1-3t+t^{2})(1+t-t^{2})},$ (213)
    $\displaystyle\sum_{k=0}^{\infty}D_{h,1}^{(k)}t^{k}=\frac{t(2-t)}{(1-3t+t^{2})(1+t-t^{2})}.$ (214)
The following table summarizes the value of $D_{f,\delta}^{(k)}$ and the conjectural value of $\dim_{\mathbb{Q}}Z_{f,\delta}^{(k)}$ for each $f\in\{g,h\}$, $\delta\in\{0,1\}$, and small weight $k\in\mathbb{Z}_{\geq{1}}$ obtained by numerical experiments.
Таблица 2. dimension evaluation for $\mathrm{I}_{g}$
$k$ 00 0 11 11 22 22 33 33 44 44 55 55
$D_{g}^{(k)}$ 11 11 22 22 44 44 88 88 1616 1616 3232 3232
$D_{g,0}^{(k)}$ 11 11 11 11 22 22 44 44 88 88 1616 1616
$D_{g,1}^{(k)}$ 00 0 11 11 22 22 44 44 88 88 1616 1616
$\dim_{\mathbb{Q}}Z_{g}^{(k)}$? 11 11 11 11 33 33 77 77 1515 1515 3131 3131
$\dim_{\mathbb{Q}}Z_{g,0}^{(k)}$? 11 11 00 0 11 11 33 33 77 77 1515 1515
$\dim_{\mathbb{Q}}Z_{g,1}^{(k)}$? 00 0 11 11 22 22 44 44 88 88 1616 1616
Таблица 3. dimension evaluation for $\mathrm{I}_{h}$
$k$ 00 0 11 11 22 22 33 33 44 44 55 55
$D_{h}^{(k)}$ 11 11 33 33 88 88 2121 2121 $5$ $5$ 144144 144144
$D_{h,0}^{(k)}$ 11 11 11 11 55 55 99 99 3030 3030 6868 6868
$D_{h,1}^{(k)}$ 00 0 22 22 33 33 1212 1212 2525 2525 7676 7676
$\dim_{\mathbb{Q}}Z_{h}^{(k)}$? 11 11 11 11 55 55 1515 1515 4646 4646 105105 105105
$\dim_{\mathbb{Q}}Z_{h,0}^{(k)}$? 11 11 00 0 33 33 88 88 2525 2525 5353 5353
$\dim_{\mathbb{Q}}Z_{h,1}^{(k)}$? 00 0 11 11 22 22 77 77 2121 2121 5252 5252
As an observation of these tables, we can check Corollary 5.11 numerically for $k=0,\dots,5$. Also, it holds $\dim_{\mathbb{Q}}Z_{f}^{(k)}=\dim_{\mathbb{Q}}Z_{f,0}^{(k)}+\dim_{\mathbb{Q}}Z_{f,1}^{(k)}$ for each $f\in\{g,h\}$, which implies that $Z_{f}^{(k)}$ may have a direct sum decomposition
$\displaystyle Z_{f}^{(k)}=Z_{f,0}^{(k)}\oplus Z_{f,1}^{(k)}.$ (215)
We do not claim that the direct sum decomposition is true, but since Remark 5.2, the equality (215) is true if so is the period conjecture (i.e. $\mathrm{per}:\mathcal{H}_{F_{f},R_{f}}\rightarrow\mathbb{C}$ is injective).
Remark 5.12.
M. Kaneko and H. Tsumura conjectured that the $\mathbb{Q}$-linear space spanned by multiple $\widetilde{T}$-values $\widetilde{T}(k_{1},\dots,k_{d})$ would have a direct sum decomposition by the parity of $d$ in [10, p.5]. This seems to be influenced by
$\displaystyle\widetilde{T}(k_{1},\dots,k_{d})=\mathrm{I}_{g}(\omega_{2}\omega_{3}^{k_{1}-1}\cdots\omega_{2}\omega_{3}^{k_{d}-1})\in\begin{cases}Z_{g,0}^{(k)},&d\text{: even},\\ Z_{g,1}^{(k)},&d\text{: odd}.\end{cases}$ (216)
in our context, and indeed the period conjecture shows their conjecture.
5.4. Alternating multiple mixed values
The alternating multiple mixed values (AMMVs for short) are special values defined by Xu–Yan–Zhao ([16]) and further studied by Charlton ([1]). AMMVs are related to the multiple $L$-values of level 44 44 and include our periods in the form of $\mathrm{I}_{g}(\eta_{1}\cdots\eta_{k})$. In this subsection, we observe that we can give a motivic interpretation of the AMMVs by the same method in Sections 5.1 and 5.2.
Definition 5.13.
For $\mathbf{k}=(k_{1},\dots,k_{d})\in\left(\mathbb{Z}_{\geq{1}}\right)^{d}$ and $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dots,\varepsilon_{d}),\boldsymbol{\sigma}=(\sigma_{1},\dots,\sigma_{d})\in\{\pm 1\}^{d}$ with $(\epsilon_{d},\sigma_{d})\neq(1,1)$, an alternating multiple mixed value (AMMV for short) is defined by
$\displaystyle M_{\boldsymbol{\sigma}}^{\boldsymbol{\varepsilon}}(\mathbf{k}):=\sum_{0<n_{1}<\cdots<n_{d}}\prod_{j=1}^{d}\frac{(1+\varepsilon_{j}(-1)^{n_{j}})\sigma_{j}^{(2m_{j}+1-\varepsilon_{j})/4}}{n_{j}^{k_{j}}}\in\mathbb{R}.$ (217)
Remark 5.14.
In [16], the sum of the right hand side of the equation 217 was defined as $n_{1}>\cdots>n_{d}>0$, but we define as $0<n_{1}<\cdots<n_{d}$ for convenience with our setting.
As mentioned in [16], AMMVs have iterated integral representations as follows. Set
$\displaystyle\omega_{0}$ $\displaystyle:=\frac{\leavevmode\nobreak\ d\lambda}{\lambda},$ (218)
$\displaystyle\omega_{+1}^{-1}$ $\displaystyle:=\frac{2\leavevmode\nobreak\ d\lambda}{1-\lambda^{2}}=\frac{\leavevmode\nobreak\ d\lambda}{\lambda-1}-\frac{\leavevmode\nobreak\ d\lambda}{\lambda+1},$ (219)
$\displaystyle\omega_{-1}^{-1}$ $\displaystyle:=\frac{-2\leavevmode\nobreak\ d\lambda}{1+\lambda^{2}}=\sqrt{-1}\left(\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\sqrt{-1}}-\frac{\leavevmode\nobreak\ d\lambda}{\lambda+\sqrt{-1}}\right),$ (220)
$\displaystyle\omega_{+1}^{+1}$ $\displaystyle:=\frac{2\lambda\leavevmode\nobreak\ d\lambda}{1-\lambda^{2}}=\frac{\leavevmode\nobreak\ d\lambda}{\lambda-1}+\frac{\leavevmode\nobreak\ d\lambda}{\lambda+1},$ (221)
$\displaystyle\omega_{-1}^{+1}$ $\displaystyle:=\frac{-2\lambda\leavevmode\nobreak\ d\lambda}{1+\lambda^{2}}=-\frac{\leavevmode\nobreak\ d\lambda}{\lambda-\sqrt{-1}}-\frac{\leavevmode\nobreak\ d\lambda}{\lambda+\sqrt{-1}}$ (222)
and
$\displaystyle\omega^{\varepsilon_{1},\varepsilon_{2}}_{\sigma}:=\begin{cases}-\omega_{\sigma}^{\varepsilon_{1}\varepsilon_{2}},&(\sigma,\varepsilon_{1},\varepsilon_{2})=(-1,1,-1),\\ \omega_{\sigma}^{\varepsilon_{1}\varepsilon_{2}},&\text{otherwise}\end{cases}$ (223)
for $\varepsilon_{1},\varepsilon_{2},\sigma\in\{\pm 1\}$. Then, AMMV is written as
$\displaystyle M_{\boldsymbol{\sigma}}^{\boldsymbol{\varepsilon}}(\mathbf{k})=\int_{\mathrm{dch}}\omega_{\sigma_{1}\cdots\sigma_{d}}^{\varepsilon_{1}}\underbrace{\omega_{0}\cdots\omega_{0}}_{k_{1}-1}\omega_{\sigma_{2}\cdots\sigma_{d}}^{\varepsilon_{2},\varepsilon_{1}}\underbrace{\omega_{0}\cdots\omega_{0}}_{k_{2}-1}\cdots\omega_{\sigma_{d}}^{\varepsilon_{d},\varepsilon_{d-1}}\underbrace{\omega_{0}\cdots\omega_{0}}_{k_{d}-1}.$ (224)
Therefore, AMMVs can be written as $\mathbb{Q}(\sqrt{{-1}})$-linear sum of multiple $L$-values of level 44 44. In addition, each differential forms in the equation 224 is $\mathrm{Gal}(\mathbb{Q}(\sqrt{-1})/\mathbb{Q})$-invariant as same as $\theta_{j}$ in the item I. Then, we have the following proposition as an analogue of Theorem 1.2.
Proposition 5.15.
For any $\mathbf{k}=(k_{1},\dots,k_{d})\in\left(\mathbb{Z}_{\geq{1}}\right)^{d}$ and $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dots,\varepsilon_{d}),\boldsymbol{\sigma}=(\sigma_{1},\dots,\sigma_{d})\in\{\pm 1\}^{d}$ with $(\epsilon_{d},\sigma_{d})\neq(1,1)$, there exists an element $M_{\boldsymbol{\sigma}}^{\mathfrak{m},\boldsymbol{\varepsilon}}(\mathbf{k})\in\left(\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{\mathbb{Q}}}^{(k)}\otimes_{\mathbb{Q}}\mathbb{Q}(\sqrt{-1})\right)^{\mathrm{Gal}(\mathbb{Q}(\sqrt{-1})/\mathbb{Q})}$ such that
$\displaystyle\mathrm{per}(M_{\boldsymbol{\sigma}}^{\mathfrak{m},\boldsymbol{\varepsilon}}(\mathbf{k}))=M_{\boldsymbol{\sigma}}^{\boldsymbol{\varepsilon}}(\mathbf{k}).$ (225)
In particular, it holds
$\displaystyle\dim_{\mathbb{Q}}\mathrm{AMMV}^{(k)}\leq 2^{k},$ (226)
where
$\displaystyle\mathrm{AMMV}^{(k)}:=\mathrm{Span}_{\mathbb{Q}}\{M_{\boldsymbol{\sigma}}^{\boldsymbol{\varepsilon}}(\mathbf{k})\leavevmode\nobreak\ |\leavevmode\nobreak\ k_{1}+\cdots+k_{d}=k\}.$ (227)
In addition, Xu–Yan–Zhao showed
$\displaystyle\dim_{\mathbb{Q}(\sqrt{-1})}\mathrm{AMMV}^{(k)}\otimes_{\mathbb{Q}}\mathbb{Q}(\sqrt{-1})=\dim_{\mathbb{Q}(\sqrt{-1})}\mathrm{MLV}_{4}^{(k)}\otimes_{\mathbb{Q}}\mathbb{Q}(\sqrt{-1}),$ (228)
where
$\displaystyle\mathrm{MLV}_{4}^{(k)}:=\mathrm{Span}_{\mathbb{Q}}\left\{L\left(\begin{matrix}k_{1},\dots,k_{d}\\ \alpha_{1},\dots\alpha_{d}\end{matrix}\right)\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ k_{1}+\cdots+k_{d}=k,\alpha_{1},\dots\alpha_{d}\in\mu_{4}\right\}$ (231)
(see [16, p.37, Theorem 6.7]). Also, Deligne showed
$\mathrm{per}\left(\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{\mathbb{Q}}}^{(k)}\right)=\mathrm{MLV}_{4}^{(k)}$
(see [3, p.37]). By combining these two facts, we have the following corollary.
Corollary 5.16.
If the period conjecture is true (i.e. $\mathrm{per}:\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{\mathbb{Q}}}\rightarrow\mathbb{C}$ is injective),
$\displaystyle\dim_{\mathbb{Q}}\mathrm{AMMV}^{(k)}=2^{k}.$ (232)
This result is consistent with Xu–Yan–Zhao’s numerical experiment ([16, p.33, Table 1]).
Список литературы
  • [1] Steven Charlton. On the evaluations of multiple ${S}$ and ${T}$ values of the form ${S}(\overset{{}_{(-)}}{2},1,\ldots,1,\overset{{}_{(-)}}{1})$ and ${T}(\overset{{}_{(-)}}{2},1,\ldots,1,\overset{{}_{(-)}}{1})$, 2024.
  • [2] K-T. Chen. Iterated integrals of differential forms and loop space homology. Ann. of Math., 97(2):217–246, 1973.
  • [3] P. Deligne. Le groupe fondamental unipotent motivique de $\mathbb{G}_{m}\backslash\mu_{N}$ pour $N=2,3,4,6$ or 88 88. Publ. Math. Inst. Hautes Etudes Sci., 112(1):101–141, 2010.
  • [4] P. Deligne and A-B. Goncharov. Groupes fondamentaux motiviques de Tate mixte. Ann. Sci. École Norm. Sup., 38(1):1–56, 2005.
  • [5] J-I-B. Gil and J. Fresán. Multiple Zeta Values: From Numbers to Motives.
  • [6] C. Glanois. Motivic unipotent fundamental groupoid of $\mathbb{G}_{m}\backslash\mu_{N}$ for $N=2,3,4,6,8$ and Galois descents. Journal of Number Theory, 160:334–384, 2016.
  • [7] A. B. Goncharov. Multiple polylogarithms and mixed Tate motives, 2001.
  • [8] A-B. Goncharov. Galois symmetries of fundamental groupoids and noncommutative geometry. Duke Math. Journal, 128(2):209–284, 2005.
  • [9] M. Kaneko K. Ihara and D. Zagier. Derivation and double shuffle relations for multiple zeta values. Compos. Math., 142:307–338, 2006.
  • [10] M. Kaneko and H. Tsumura. Multiple ${L}$-values of level four, poly-Euler numbers, and related zeta functions. arXiv:2208.05146, 2022.
  • [11] K-S. Kedlaya. $p$-adic cohomology: from theory to practice. In $p$-adic geometry, Univ. Lecture Ser., 45, pages 175–203. Amer. Math. Soc., 2008.
  • [12] T-Q-T. Le and J. Murakami. Representations of the category of tangles by Kontsennch’s itemted integral. Commun. Math. Phys., 168:535–562, 1995.
  • [13] Y. Ohno. A generalization of the duality and sum formulas on the multiple zeta values. J. Number Theory, 74:39–43, 1999.
  • [14] E. Otsuka. On arithmetic properties of periods for some rational differential forms over $\mathbb{Q}$ on the Fermat curve $F_{2}$ of degree 2. arXiv:2302.07650, 2023.
  • [15] T. Terasoma. Mixed Tate motives and multiple zeta values. Invent. Math., 149(2):339–369, 1001.
  • [16] Ce Xu, Lu Yan, and Jianqiang Zhao. Alternating Multiple Mixed Values, 2023.
  • [17] D. Zagier. Values of zeta functions and their applications. In First European Congress of Mathematics Paris, July 6–10, 1992: Vol. II: Invited Lectures (Part 2), pages 497–512. Springer, 1994.